### v1

# Convergence of the open WASEP stationary measure without Liggett’s condition

Abstract.
We demonstrate that it is not necessary to assume Liggett’s condition in order to obtain convergence of the open ASEP stationary measures to the open KPZ stationary measure. This is equivalent to demonstrating that, under weak asymmetry scaling and appropriate scaling of time and space, the four-parameter Askey-Wilson process converges to a two-parameter continuous dual Hahn process. We conjecture that the convergence of the open ASEP height function process to solutions to the open KPZ equation will hold for a wider range of ASEP parameters than those permitted by Liggett’s condition.
1. Introduction
1.1. Preface
The open asymmetric simple exclusion process (ASEP), an interacting particle system of great interest in probability and mathematical physics, first appeared in biology, in a 1968 work of MacDonald, Gibbs, and Pipkin [MGP68] studying the kinetics of protein synthesis. In subsequent years, it also appeared in an early study of exclusion processes by Spitzer [Spi70] and Liggett [Lig75]. The open ASEP is a Markov process which models particles moving on a one-dimensional lattice of finite size with open boundaries, allowing particles to enter and exit the lattice on either side. The process takes its name from its properties: it is open in the sense that particles may enter and exit at either boundary, the left and right rates of particle movement are asymmetric, meaning that particles move to the right at a higher rate than that at which they move to the left; and only one particle may occupy any location at a given time, a property called exclusion. There are other, well studied variants of this system, including a similar model on an infinite lattice, a half infinite lattice, or a periodic domain.
The open ASEP is a relatively simple model of a system with nonzero net current in its equilibrium state, and, despite its simple definition, it exhibits complex behaviors associated to many-particle systems, such as phase transitions. Due to these properties, the open ASEP has been of great interest to researchers in probability and mathematical physics. Bertini and Giacomin [BG97] discovered that the ASEP on the full line converges to the Kardar-Parisi-Zhang (KPZ) equation on the full line under weak asymmetry scaling assumptions (see Section 1.3 for a longer discussion of this relationship). Corwin and Shen [CS18] then discovered that the Open ASEP on two additional domains (half-open and open ASEP) converge to the half-space and open KPZ equations, respectively, under triple point scaling of the boundary parameters and weak asymmetry scaling. Parekh [Par19] extended the results from Corwin and Shen which were restricted to a certain range of parameters of ASEP to hold for all parameters. As discovered by Bertini and Giacomin [BG97], the open ASEP is closely related to the Kardar-Parisi-Zhang (KPZ) equation (see Section 1.3 for a longer discussion of this relationship). Corwin and Knizel [CK24] used the connection between the open ASEP and the open KPZ equation to give the first construction of the stationary measure for the open KPZ equation through a scaling limit of open ASEP stationary measures. Work by Derrida, Evans, Hakim, and Pasquier [DEHP93] introduced an algebraic technique, called the matrix product ansatz (MPA) for obtaining the stationary measure of the open ASEP. Uchiyama, Sasamoto, and Wadati [USW04] used the MPA expression for the stationary measure of open ASEP to express the measure in terms of Askey-Wilson polynomials. More recently, Bryc and Wesolowski [BW17] demonstrated a formula for the multi-point correlation function of the stationary measure in terms of a process called the Askey-Wilson Process (which has one-point and transition densities given in terms of Askey-Wilson polynomials). Corwin and Knizel characterized the Laplace transform of the finite dimensional marginals of the stationary distribution of the open KPZ equation in terms of a stochastic process called the continuous dual Hahn process (see Section 2.3), which is the limit of the Askey-Wilson process under the appropriate scaling.
The KPZ equation has a unique stationary measure, which was proved by Knizel and Matetski [KM22] for $u+v>0$ and $\min(u,v)>-1$, and by Parekh for all $u,v\in\mathbb{R}$. Corwin and Knizel characterized this stationary measure in terms of its Laplace transform for all $u+v\geq 0$. Barraquand and Le Doussal [BLD22] and Bryc, Kuznetsov, Wang, and Wesolowski [BKWW23] showed that the stationary measure of the open KPZ equation constructed by Corwin and Knizel, has a probabilistic description as a process $\{H_{u,v}(x)\}_{x\in[0,1]}=\{\mathbb{B}(\frac{x}{2})-\mathbb{X}_{u,v}(x)\}_{x\in[0,1]}$, where $\mathbb{B}(x)$ is standard Brownian motion and $\mathbb{X}_{u,v}(x)$ is an exponential re-weighting of Brownian motion by a Radon-Nikodym derivative
[…]
Where $f\in C[0,1]$. For the purpose of this paper, we will use the description of the open ASEP in terms of its multipoint Laplace transform, see (3.2). Recently, work by Barraquand, Corwin, and Yang [BCY] constructed the KPZ stationary measure for all $u,v\in\mathbb{R}$ as a scaling limit from the log-gamma directed polymer model, modulo a result extending the existing understanding of convergence of the spacetime process of the log-gamma polymer to convergence on an interval $[0,L]$ (see Claim 4.2, [BCY]).
Corwin and Knizel’s construction of the open KPZ stationary measure assumed a relation, called Liggett’s condition, on the ASEP parameters. This relation was also used in earlier works by Corwin and Shen [CS18] and Parekh [Par19] that showed that the open ASEP height function process converges as a space-time process to solutions of the open KPZ. The contribution of this paper is to demonstrate that Liggett’s condition is not necessary to obtain convergence of the stationary measures of open ASEP to that of open KPZ. Based on this result, we conjecture that the convergence of open ASEP to open KPZ as a space-time process will also hold without this assumption.
This paper also demonstrates an interesting property of the Askey-Wilson process and its limiting process, the continuous dual Hahn process. The Askey-Wilson process is defined in general with parameters $(A,B,C,D,q)$ (see 2.1), but Liggett’s condition fixes the values of $B$ and $D$ as functions of $q$. The fact that the limit of the Askey-Wilson process when $B$ and $D$ are allowed to vary is still a continuous dual Hahn process which depends only on $A$ and $C$ is an interesting observation about the structure of the Askey-Wilson and continuous dual Hahn processes, which have been studied in works such as [BW10, BWW23, BW19, Bry22, WWY]. This observation was anticipated by Bryc, Wang, and Wesolowski [BWW23] (see, for example, the remarks before Assumption 1.1 in this paper).
1.2. The Open ASEP
To better understand Liggett’s condition, we begin by giving a more detailed introduction of the open ASEP.
[…]
Figure 1. The open ASEP lattice with right, left, and boundary jump rates
The underlying lattice of the open ASEP has finite size $N$. We use $\alpha,\gamma$ to denote the rate of particle movement into and out of the leftmost lattice point from, respectively to, the left reservoir. Similarly, we use $\beta,\delta$ to denote the rate of particle movement out of, respectively into, the rightmost lattice point of the lattice to, respectively from, the right reservoir. Within the lattice, particles jump at times which are determined by independent, exponentially distributed random variables, with rate $1$ for movement to the right and rate $q\in(0,1)$ for movement to the left. The state space of the open ASEP is $\{0,1\}^{N}$. We refer to elements of the state space by $\tau=(\tau_{1}(t),...,\tau_{N}(t)$ where $\tau_{i}(t)=1$ if and only if there is a particle in position $i$ at time $t$. The open ASEP process $\tau(t)$ is a Markov process on the state space $\{0,1\}^{N}$ and the infinitesimal generator $\mathcal{L}(\cdot)$ which acts on functions of the open ASEP state space $f:\{0,1\}^{N}\to\mathbb{R}$ by
$(\mathcal{L}f)(\tau)$ $=\left(\alpha(1-\tau_{1})+\gamma\tau_{1}\right)\cdot\left(f(\tau^{1})-f(\tau)\right),$
$+\sum_{i=1}^{N-1}\left(\tau_{i}(1-\tau_{i+1})+q(1-\tau_{i})\tau_{i+1}\right)\cdot\left(f(\tau^{i,i+1})-f(\tau)\right),$
$+\left(\beta\tau_{N}+\delta(1-\tau_{N})\right)\cdot\left(f(\tau^{N})-f(\tau)\right).$
In this equation, the notation $\tau^{i}$ means that, in $\tau$, the value of $\tau_{i}$ is sent to $1-\tau_{i}$ and the notation $\tau^{i,i+1}$ means that $\tau_{i}$ and $\tau_{i+1}$ are swapped. We will denote the stationary measure of the open ASEP with parameters $\alpha,\beta,\gamma,\delta,q$ and lattice size $N$ as $\pi_{N}(\cdot)$, suppressing the dependence on the boundary parameters and the asymmetric jump rate. For the duration of this work, we will use $\langle f\rangle_{N}$ to denote the expectation of a function $f$ under $\pi_{N}(\cdot)$. The time evolution of the open ASEP induces a Markov process, called the height function process,
$h_{N;\alpha,\beta,\gamma,\delta}(t,x)=h_{N;\alpha,\beta,\gamma,\delta}(t,0)+\sum_{i=1}^{\lfloor Nx\rfloor}(2\tau_{i}(t)-1),$
where $h_{N;\alpha,\beta,\gamma,\delta}(t,0)$ is the net number of particles which have entered through the left boundary at time $t$. If $\tau$ is distributed according to $\pi_{N}(\cdot)$, we refer to the law of the process $x\mapsto h_{N;\alpha,\beta,\gamma,\delta}(t,x)-h_{N;\alpha,\beta,\gamma,\delta}(t,0)$ as the stationary measure for the open ASEP height function increment process.
One other standard quantity associated to ASEP is the stationary current, or the average net current of particles entering the ASEP system. As the size of the lattice grows, the value of the limiting stationary current depends on two parameters, $\rho_{\ell}$ and $\rho_{r}$, which are effective densities at the left and right boundaries. Changing parameters to $(A,B,C,D,q)$,
$A:=\kappa^{+}(q,\beta,\delta),$ $B:=\kappa^{-}(q,\beta,\delta),$ $C:=\kappa^{+}(q,\alpha,\gamma),$ $D:=\kappa^{-}(q,\alpha,\gamma),$
$\kappa^{\pm}(q,x,y)=\frac{1}{2x}\left(1-q-x+y\pm\sqrt{(1-q-x+y)^{2}+4xy}\right),$
we can express $\rho_{\ell}$ and $\rho_{r}$ as
$\rho_{\ell}=\frac{1}{1+C},$ $\rho_{r}=\frac{A}{1+A}.$
We note that $\rho_{\ell}$ and $\rho_{r}$ are independent of the parameters $B$ and $D$. For a thorough description of the dynamics of open ASEP in terms of parameters $\rho_{\ell}$ and $\rho_{r}$, see the survey [Cor22].
The open ASEP stationary measure is known to converge to the open KPZ stationary measure as $N\to\infty$ and under the scaling
(1.1) $A=q^{v},$ $B=-q,$ $C=q^{u},$ $D=-q,$ $q=\exp{\left(\frac{-2}{\sqrt{N}}\right)},$
and a particular choice of scaling for the height function (see 1.4). This choice of scaling for $q$ is called weak asymmetry scaling. Corwin and Knizel showed that for any $u,v\in\mathbb{R}$, the $N$-indexed collection of open ASEP stationary measures, scaled according to the weak asymmetry scaling in $N$, was tight, and that sub-sequential limits were stationary measures of the open KPZ equation. Under the constraint that $u+v>0$, they also showed that all sub-sequential limits were equal, and gave a formula for the stationary measures in terms of their Laplace transforms. Later results by Knizel and Matetski [KM22] (for $u+v>0$ and $\min\{u,v\}>-1$) and and Parekh [Par] (for all $u,v\in\mathbb{R}$) demonstrated that the open KPZ stationary measure is unique.
The scaling of $A,B,C,D$ with respect to $q$ was chosen to preserve two properties. It preserves a $N^{-\frac{1}{2}}$ scaling around the triple point,
$\rho_{\ell}=\frac{1}{2}+\frac{u}{2}N^{-\frac{1}{2}}+o(N^{-\frac{1}{2}}),$ $\rho_{r}=\frac{1}{2}-\frac{v}{2}N^{-\frac{1}{2}}+o(N^{-\frac{1}{2}}),$
by requiring $A=q^{v}$ and $C=q^{u}$ for some $u,v\in\mathbb{R}$. It also preserves a property called Liggett’s condition. Liggett’s condition is a constraint on the boundary parameters of the system, which, in the parameterization $(A,B,C,D)$, implies that $B=D=-q$. It is typically stated in terms of $(\alpha,\beta,\gamma,\delta)$ as
(1.2) $\alpha+\gamma/q=1,$ $\beta+\delta/q=1.$
The contribution of this paper is to observe that, even without assuming Liggett’s condition, the scaling limit of the stationary measure of the open ASEP still converges to the stationary measure of the open KPZ equation. The new scaling assumptions are collected below.
Assumptions 1.1.
Our scaling assumptions are
  1. (1)
    Weak asymmetry scaling: $q=\exp{\left(\frac{-2}{\sqrt{N}}\right)}.$
  2. (2)
    Scaling of boundary parameters
    (1.3) $A=q^{v},$ $B=-q^{w},$ $C=q^{u},$ $D=-q^{r},$
    where $u+v>0$ and $w,r>0$.
  3. (3)
    Height function scaling: For $h_{N}(\cdot)$ the height function of the open ASEP, and $x\in[0,1]$ we define
    $h^{(N)}_{u,v,w,r}(x)$ $:=N^{-\frac{1}{2}}h_{N;\alpha,\beta,\gamma,\delta}\left(Nx\right).$
    We will often drop the subscript and simply use the notation $h^{(N)}(x).$
Condition $(2)$ is interchangeable with the following condition on boundary rates $(\alpha,\beta,\gamma,\delta)$,
(1.4) $\alpha=\frac{1-q}{(1-q^{r})(1+q^{u})},$ $\beta=\frac{1-q}{(1-q^{w})(1+q^{v})},$
$\gamma=\frac{q^{u+r}(1-q)}{(1-q^{r})(1+q^{u})},$ $\delta=\frac{q^{v+w}(1-q)}{(1-q^{w})(1+q^{v})}.$
To demonstrate that the wider range of parameters captures a wider range of open ASEP stationary measures, we consider the case where $N=2$ and then separately compute the stationary measure in the case where $q=\frac{1}{2}$ and $u=v=1$, with $w=r=1$ and $w=2,r=3$.
Example 1.2.
Fixing $\rho_{\ell}=(1+q)^{-1},\rho_{r}=q(1+q)^{-1}$ (equivalently, $u=v=1$), we vary $w$ and $r$:
  1. (1)
    Setting $w=r=1$, we have $(\alpha,\beta,\gamma,\delta)=\left(\frac{2}{3},\frac{2}{3},\frac{1}{6},\frac{1}{6}\right)$. The stationary measure is given by
    ($\mathbb{P}(\{0,0\})=\frac{1}{4},$ ($\mathbb{P}(\{1,0\})=\frac{1}{4},$ ($\mathbb{P}(\{0,1\})=\frac{1}{4},$ ($\mathbb{P}(\{1,1\})=\frac{1}{4}.$
  2. (2)
    Setting $w=2,r=3$, we obtain $(\alpha,\beta,\gamma,\delta)=\left(\frac{8}{21},\frac{4}{9},\frac{1}{14},\frac{1}{18}\right)$ and stationary measure is given by
    ($\mathbb{P}(\{0,0\})$ $=\frac{35845251}{140695480}\approx 0.255$ ($\mathbb{P}(\{1,0\})=\frac{7259395}{28139096}\approx 0.258$
    ($\mathbb{P}(\{0,1\})$ $=\frac{6549543}{28139096}\approx 0.233$ ($\mathbb{P}(\{1,1\})=\frac{35805539}{140695480}\approx 0.254.$
This example illustrates that altering the parameters $w$ and $r$ effects the stationary measure of the open ASEP when $N$ is finite. The result of this paper demonstrates that these small differences disappear in the scaling limit, and the limiting stationary measure depends only on $u$ and $v$. It would be interesting to study how the rate of convergence of the open ASEP stationary measures varies in $w$ and $r$.
1.3. Statement and Discussion of the Theorem
The open KPZ equation is a stochastic partial differential equation defined for $t\in\mathbb{R}_{\geq 0}$ and $x\in[0,1]$ by
$\partial_{t}H(t,x)=\frac{1}{2}\partial^{2}_{x}H(t,x)+\frac{1}{2}\left(\partial_{x}H(t,x)\right)^{2}+\xi(t,x),$
where $\xi(t,x)$ is space-time white noise. Solutions of the open KPZ equation satisfy Neumann boundary conditions $(\partial_{x}H)(t,0)=u$ and $(\partial_{x}H)(t,1)=-v$. This definition is slightly informal. For background information on the open KPZ equation and a careful definition, we direct the reader to Corwin and Shen [CS18] or Parekh [Par19]. Both of those works proved convergence of the open ASEP height function process to Hopf-Cole solutions to the open KPZ equation for a range of boundary parameters under the scaling conditions in 1.4, below.
The unique open KPZ stationary measure is the law of the process $x\mapsto H_{0}(x)-H_{0}(0)$ on $(C[0,1],\mathcal{B}(C[0,1]))$ such that the law of the space-time process $(t,x)\mapsto H(t,x)-H(t,0)$, produced by evolving initial data $H(0,x)-H(0,0)$ (with $H_{0}(x)=H(0,x)$) according to the open KPZ equation, is invariant in time. Corwin and Knizel [CK24] gave an explicit formula (for $u+v\geq 0$) for the open KPZ stationary measure in terms of the Laplace transform of the continuous dual Hahn process (3.2). We will denote the stationary measure of the open KPZ equation with boundary parameters $u,v\in\mathbb{R}$ as the law of the process $x\mapsto H_{u,v}(x)-H_{u,v}(0)$ for $x\in[0,1]$. Using this notation, we can state the main result of this article.
Theorem 1.3.
Assume that the open ASEP parameters and height function satisfy the scaling parameters in 1.1. The scaling induces a tight sequence of laws of $\{h_{u,v,w,r}^{(N)}(\cdot)-h_{u,v,w,r}^{(N)}(0)\}_{N\in\mathbb{N}}$ on $(C[0,1],\mathcal{B}(C[0,1]))$ and for all $u+v>0$ and $w,r>0$, this sequence converges weakly to the law of $H_{u,v}(\cdot)-H_{u,v}(0)$, the stationary measure of the open KPZ equation, a limit which does not depend on $w,r$.
It is natural to consider which of these scaling assumptions is necessary to obtain the convergence of the open ASEP space-time process to that of the open KPZ. Convergence results for various ASEP models to KPZ variants by Corwin and Shen [CS18]; Parekh [Par19]; and Yang [Yan22] have followed the framework of a foundational result by Bertini and Giacomin [BG97], who approached the problem via the Hopf-Cole transform. Consequently, the argument in each of these works requires Liggett’s condition. While Gonçalves, Perkowski and Simon[GPS20] do not use the Hopf-Cole transform, they still assume Liggett’s condition. It is not known whether convergence as processes occurs under the full range of boundary parameters for the open ASEP. The effective boundary densities, $\rho_{\ell}$ and $\rho_{r}$, depend only on $A$ and $C$, which suggests we may vary $B$ and $D$ without disrupting the triple point scaling. Our result in 1.3 provides further evidence that it is possible to obtain convergence as space-time processes without Liggett’s condition, since it is possible at the level of stationary measures. We conjecture that convergence as space-time processes holds using the scaling assumptions in 1.1.
Assumptions 1.4.
In [CS18], the scaling assumptions for the convergence of the open ASEP height function $h_{N}(\cdot)$ to the Hopf-Cole solution to the open KPZ equation are
  1. (1)
    Weak asymmetry scaling: $q=\exp{\left(\frac{-2}{\sqrt{N}}\right)}$
  2. (2)
    Liggett’s condition:
    $\alpha+\gamma/q=1$ $\beta+\delta/q=1$
  3. (3)
    4:2:1 Height function scaling: For $h_{N}(\cdot)$ the height function process of the open ASEP, $t\geq 0$ and $x\in[0,1]$ we define
    $h^{(N)}_{u,v}(t,x)$ $:=N^{-\frac{1}{2}}h_{N;\alpha,\beta,\gamma,\delta}\left(\frac{1}{2}e^{N^{-\frac{1}{2}}}N^{2}t,Nx\right)+\left(\frac{1}{2}N^{-1}+\frac{1}{24}\right)t$
    $z^{(N)}_{u,v}(t,x)$ $:=\exp{\left(h_{u,v}^{(N)}(t,x)\right)}$
  4. (4)
    Hölder bounds on initial data: The open ASEP initial data is $N$-indexed and satisfies Hölder bounds. For all $\theta\in\left(0,\frac{1}{2}\right)$ and $n\in\mathbb{Z}_{n\geq 1}$, there exist constants $C(n)$ and $C(\theta,n)>0$ such that for all $x,x^{\prime}\in[0,1]$ and all $N\in\mathbb{Z}_{\geq 1}$,
    $\mathbb{E}\left[|z^{(N)}_{u,v}(x)|^{n}\right]^{1/n}\leq C(n)$ $\mathbb{E}\left[|z^{(N)}_{u,v}(x)-z^{(N)}_{u,v}(x^{\prime})|^{n}\right]^{1/n}\leq C(\theta,n)|x-x^{\prime}|^{\theta}$
    where $\mathbb{E}[\cdot]$ is the expectation over ASEP height function initial data $h_{N}(x)$.
Conditions $(1)$ and $(3)$ are the same as the equivalently numbered conditions in 1.1 in the stationary case, but Liggett’s condition is an additional assumption.
Conjecture 1.5.
We denote the Hopf-Cole solution to the open KPZ equation with Neumann boundary parameters $u$ and $v$ by $H(t,x)$. We define $Z(t,x):=\exp{\left(H(t,x)\right)}$. We denote the diffusively scaled height function process with boundary rates $\alpha,\beta,\gamma,\delta$ given by (1.4) as $h_{u,v,w,r}^{(N)}(t,x)$, and $z_{u,v,w,r}^{(N)}(t,x):=\exp{\left(h_{u,v,w,r}^{(N)}(t,x)\right)}$. Under the scaling assumptions in 1.1, and using the height function process scaling defined in 1.4 (4) and requiring the same Hölder bounds described in 1.4 (5), then, if the initial data $z^{(N)}_{u,v,w,r}(0,x)$ converges to $Z(0,x)$, $z^{(N)}_{u,v,w,r}(t,x)\implies Z(t,x)$ as a space-time process.
1.4. Approach to proving 1.3
The proof of the tightness result in 1.3 is an application of an attractive coupling of ASEP proved in [CK24]. This result allows us to obtain Hölder bounds on $h^{(N)}_{u,v,w,r}(x)-h^{(N)}_{u,v,w,r}(0)$ which consequently allow us to conclude that the sequence is tight. For the details of this approach, see Section 5
In order to prove that the sequence of measures in 1.3 converges to the law of $H_{u,v}(\cdot)-H_{u,v}(0),$ we rely on the fact that the open ASEP stationary measure can be written in terms of a stochastic process called the Askey-Wilson process. Consequently, we can describe the Laplace transform of the height function process in terms of the Askey-Wilson process and analyze the asymptotics of that expression in terms of Askey-Wilson densities. We also rely on the fact that pointwise convergence of Laplace transforms on an open set implies convergence in distribution (see 2.6). The main technical contribution of this paper is to show that in a more general range of parameters, for any $w,r>0$ and $u+v>0$, the Askey-Wilson process in parameters $v,w,u,r$ (see Section 2.2.2) converges to the continuous dual Hahn process with parameters $u,v$. Consequently, when $u+v>0$ and $w,r>0$, the tight sequence of laws from 1.3 has a unique limit, which is the stationary measure of the open KPZ equation.
1.5. Organization of the Paper
In Section 2, we introduce the Askey-Wilson process and the continuous dual Hahn process, and state results and definitions pertaining to the asymptotics of q-pochammer symbols. In Section 3, we state the main technical lemma, 3.1, that the Laplace transforms of the open ASEP height functions $\{h_{u,v,w,r}^{(N)}(\cdot)-h_{u,v,w,r}^{(N)}(0)\}_{N\in\mathbb{N}}$ converge pointwise to the Laplace transform of the open KPZ stationary measure when $u+v>0$ and $w,r>0$. The remaining parts of Section 3 demonstrate that the scaling limit of the Askey-Wilson measures produces the continuous dual Hahn process under the scaling assumptions in 1.1. In Section 4 we prove 3.1. In Section 5, we combine these results to prove 1.3.
1.6. Acknowledgements
The author thanks Ivan Corwin for suggesting the problem, and Ivan Corwin and Alisa Knizel for helpful discussions about their work. The author also thanks Włodek Bryc and Yizao Wang for helpful comments on a draft. The author’s research was supported by Ivan Corwin’s NSF grant DMS-1811143 and the Fernholz Foundation’s “Summer Minerva Fellows” program
2. Background
In Section 2.2 and Section 2.3 we will characterize the Askey-Wilson process and the continuous dual Hahn process, which are used to describe the stationary measures of the open ASEP and open KPZ, respectively. In Section 2.4 we will quote asymptotic results on q-pochammer symbols, proved in [CK24], which will be necessary for our argument. In Section 2.5, we formally state the result that obtains weak convergence of random variables from pointwise convergence of their Laplace transforms.
2.1. General Notation
Throughout the article, we will use the notation $(a;q)_{k}:=\prod_{j=0}^{k-1}(1-aq^{j})$ for $q$-Pochhammer symbols, and the shorthand $(a_{1},...,a_{n};q)_{k}:=(a_{1},q)_{k}\cdot\cdot\cdot(a_{n};q)_{k}$. We will also often use $(a)_{k}$ and $(a_{1},...,a_{n})_{k}$ to refer to the same quantities, removing $q$ from the notation, since for the entire paper we will set $q=\exp{\left(\frac{-2}{\sqrt{N}}\right)}$. We will likewise use $[a]_{k}:=\prod_{j=0}^{k-1}(a+j)$ to denote the standard Pochhammer symbol and a similar shorthand for products, $[a_{1},...,a_{n}]_{k}:=[a_{1}]_{k}\cdot\cdot\cdot[a_{n}]_{k}$. We will use the notation $\Gamma(x_{1},...,x_{n}):=\Gamma(x_{1})\cdot\cdot\cdot\Gamma(x_{n})$ to denote a product of Gamma functions.
2.2. The Askey-Wilson Process
The stationary measure for the open ASEP can be described in terms of a stochastic process called the Askey-Wilson process. This process first appeared in the work of Bryc and Wesolowski [BW17], who formulated the probability generating function of the open ASEP in terms of the Askey-Wilson process $\mathbb{Y}_{t}$, see 2.1,
$\left\langle\prod_{j=1}^{N}t_{j}^{\tau_{j}}\right\rangle_{N}=\frac{\mathbb{E}\left[\prod_{j=1}^{N}(1+t_{j}+2\sqrt{t_{j}}\mathbb{Y}_{t_{j}})\right]}{2^{N}\mathbb{E}\left[(1+\mathbb{Y}_{1})^{N}\right]}.$
In this paper, we will use the Askey-Wilson formula to characterize the Laplace transform of the open ASEP height function. We recall the definition of the diffusively scaled height function, $h^{(N)}(x)$ from 1.1. We can write its multipoint Laplace transform as
(2.1) $\phi^{(N)}_{u,v,w,r}(\textbf{c},\textbf{x})$ $=\left\langle e^{-\sum_{k=1}^{d}c_{k}\left(h^{(N)}_{u,v,w,r}(x_{k}^{(N)})-h^{(N)}_{u,v,w,r}(0)\right)}\right\rangle_{N}$
$=\frac{\mathbb{E}\left[\prod_{k=1}^{d+1}\left(\cosh(s_{k}^{(N)})+\mathbb{Y}_{e^{-2s_{k}^{(N)}}}\right)^{n_{k}-n_{k-1}}\right]}{\mathbb{E}\left[(1+\mathbb{Y}_{1})^{N}\right]}.$
In this expression, as in the rest of the paper, we use the notation
(2.2) x $:=(x_{0},...,x_{d+1}),\text{ where }0=x_{0}<x_{1}<...<x_{d}\leq x_{d+1}=1,$
c $:=(c_{1},...,c_{d}),\text{ where }c_{1},...,c_{d}>0,$
$s_{k}$ $=c_{k}+...+c_{d},\text{ and }s_{d+1}=0.$
We will also use the notation $n_{k}=\lfloor Nx_{k}\rfloor$, $s_{k}^{(N)}=N^{-\frac{1}{2}}(s_{k}+c)$ and $x_{k}^{(N)}=N^{-1}\lfloor Nx_{k}\rfloor$. The rest of this subsection is devoted to defining the Askey-Wilson process.
Definition 2.1 (Section 1.3 [BW17]).
The Askey-Wilson process $\{\mathbb{Y}_{t}\}_{t\geq 0}$ is the Markov process whose marginal and transitional distributions on $V\in\mathcal{B}(\mathbb{R})$ are given by
(2.3) $\pi_{t}(V)$ $:=AW\left(V;A\sqrt{t},B\sqrt{t},\frac{C}{\sqrt{t}},\frac{D}{\sqrt{t}}\right),$
$\pi_{s,t}(x,V)$ $:=AW\left(V;A\sqrt{t},B\sqrt{t},\sqrt{\frac{s}{t}}(x+\sqrt{x^{2}-1}),\sqrt{\frac{s}{t}}(x-\sqrt{x^{2}-1})\right).$
In these expressions, the functions $AW(V;a,b,c,d)$ are the Askey-Wilson measures. These measures have a continuous support and a discrete support, which we denote $S^{c}(a,b,c,d,q)$ and $S^{d}(a,b,c,d,q)$, respectively. This measure is given by
[…]
$AW^{c}(x;a,b,c,d,q)$ $:=\mathbf{1}_{|x|<1}\frac{(q,ab,ac,ad,bc,bd,cd)_{\infty}}{2\pi(abcd)_{\infty}\sqrt{1-x^{2}}}\bigg{|}\frac{(e^{2i\theta_{x}})_{\infty}}{(ae^{i\theta_{x}},be^{i\theta_{x}},ce^{i\theta_{x}},de^{i\theta_{x}})_{\infty}}\bigg{|}^{2}$
$AW^{d}(V;a,b,c,d,q)$ $:=\sum_{y\in V\cap S^{d}(a,b,c,d,q)}AW^{d}(y;a,b,c,d,q).$
We refer to $AW^{c}(x;a,b,c,d,q)$ as the continuous density at $x$ and $AW^{d}(y;a,b,c,d,q)$ as the discrete mass at $y$. When we allow the parameter $q$ to scale in $N$, we will denote the corresponding one-point and transition densities by $\pi_{s}^{(N)}(\cdot)$ and $\pi_{s,t}^{(N)}(\cdot,\cdot)$.
The discrete support of the Askey-Wilson measure depends on the parameters $\chi\in\{a,b,c,d\}$. Atoms exist whenever one of these parameters satisfies $|\chi|>1$. If one or more of the parameters satisfies this condition, then each $\chi$ with $|\chi|>1$ will have an associated collection of atoms, which we will denote by $\{y_{j}^{\chi}\}_{j\in J}$. The index set $J$ consists of all $j$ such that $|q^{j}\chi|>1$. The atoms $y_{j}^{\chi}$ will have locations given by $y_{j}^{\chi}:=\frac{1}{2}\left(q^{j}\chi+\frac{1}{q^{j}\chi}\right).$ In the case where $\chi=a$, the discrete mass at each of the $y_{j}^{a}$ is given by
(2.4) $AW^{d}(y_{0}^{a};a,b,c,d,q)$ $:=\frac{(a^{-2},bc,bd,cd)_{\infty}}{(b/a,c/a,d/a,abcd)_{\infty}},$
$AW^{d}(y_{j}^{a};a,b,c,d,q)$ $:=AW^{d}(y_{0}^{a};a,b,c,d,q)\frac{(a^{2},ab,ac,ad)_{j}(1-a^{2}q^{2j})q^{j}}{(q,qa/b,qa/c,qa/d)_{j}(1-a^{2})(abcd)^{j}}.$
When $\chi\neq a$, the expressions for the weights are the same, with $\chi$ and $a$ swapped throughout the expression. We use the notation $y_{j}^{\chi}(t)$ to represent an atom of the measure $\pi_{t}(V)$.
2.2.1. One Point Distribution Atoms
In 2.1, the one-point distribution $\pi_{t}(V)$ takes the form of an Askey-Wilson measure $AW(V;a,b,c,d,q)$ with parameters
$a=A\sqrt{t}=q^{v}\sqrt{t},$ $b=B\sqrt{t}=-q^{w}\sqrt{t},$ $c=C/\sqrt{t}=q^{u}/\sqrt{t},$ $d=D/\sqrt{t}=-q^{r}/\sqrt{t}.$
In order to simplify the asymptotic computations, we will deal with a rescaled process, which is indexed by time variable $q^{t}$, so that
$a=q^{v+\frac{t}{2}},$ $b=-q^{w+\frac{t}{2}},$ $c=q^{u-\frac{t}{2}},$ $d=-q^{r-\frac{t}{2}}.$
Atoms of the Askey-Wilson measure arise whenever $\chi\in\{a,b,c,d\}$ has the property that $|\chi|>1$. This can occur in one of four cases:
[…]
In the work of Corwin and Knizel [CK24], which we recover by setting $w=r=1$, the domain in $t$ on which convergence occurs is restricted to an interval $(-2,2)$ in order to prevent atoms from arising from cases $(2)$ and $(4)$. When $w$ and $r$ are allowed to vary, there is a comparable domain in which there are no atoms arising from cases $(2)$ and $(4)$. For any $t\in(-2w,2r)$ we will avoid all such atoms. In particular, since $w,r>0$, we note that $w+r>0$, and therefore $(-2w,2r)$ is an open interval. As a reminder for why this is necessary, note that establishing convergence in distribution through the Laplace transform, as described in 2.6, requires convergence on an open set.
Likewise, atoms exist in cases $(1)$ and $(3)$ when $v$ and $u$ are respectively less than $1$. Since we constrain $AC=q^{u+v}<1$, we know that $u+v>0$ and therefore only one of the conditions $(1)$ or $(3)$ can hold at any given time. This fact will be important to calculations in Section 3.
2.2.2. Atoms and Time Reversal
In the discussion of the asymptotics of the transition probabilities, we center and rescale the Askey-Wilson Process. If the original Askey-Wilson process is $\mathbb{Y}$, characterized by $q,A,B,C,D,$ then we define the tangent process $\hat{\mathbb{Y}}^{(N)}_{s}=2N(1-\mathbb{Y}_{q^{s}})$. This parameterization involves a time reversal of $\mathbb{Y}$, and the transition probabilities of this process are related to the transition probabilities of the original process by a conjugation, which we will describe below.
We denote the measure associated to $\hat{\mathbb{Y}}_{s}$ by $\hat{\pi}_{s}^{(N)}$. As before, this measure has a discrete support and a continuous support, given by
$\hat{S}_{s}^{(N),c}=\hat{S}^{(N),c}:=[0,4N]$ $\hat{S}_{s}^{(N),d}:=\left\{y\in\mathbb{R}:1-\frac{y}{2N}\in S^{d}_{q^{s}}\right\}.$
In the continuous component of the support, we will additionally emphasize the dependence on the parameters $w,v,u,r$ by writing $\hat{\pi}_{s;w,v,u,r}^{(N)}$. In this notation, we can define $\hat{\pi}_{t}$ and $\hat{\pi}_{s,t}$ in terms of $\pi_{t}$ and $\pi_{s,t}$ via the following transformations.
$\hat{\pi}_{t;v,w,u,r}^{(N),c}(y)$ $:=\frac{1}{2N}\pi_{q^{t}}^{(N),c}\left(1-\frac{y}{2N}\right),$ $y\in\hat{S}_{t}^{(N),c},$
$\hat{\pi}_{t}^{(N),d}(y)$ $:=\pi_{q^{t}}^{(N),d}\left(1-\frac{y}{2N}\right),$ $y\in\hat{S}^{(N),d}_{t}.$
$\hat{\pi}_{s,t;v,w,u,r}^{(N),c,c}(x,y)$ $:=\frac{1}{2N}\pi_{q^{t},q^{s}}^{(N),c,c}\left(1-\frac{y}{2N},1-\frac{x}{2N}\right)\frac{\pi_{q^{t}}^{(N),c}(1-\frac{y}{2N})}{\pi_{q^{s}}^{(N),c}(1-\frac{x}{2N})},$ $x\in\hat{S}_{s;v,w,u,r}^{(N),c},y\in\hat{S}_{t}^{(N),c},$
$\hat{\pi}_{s,t;v,w,u,r}^{(N),d,c}(x,y)$ $:=\frac{1}{2N}\pi_{q^{t},q^{s}}^{(N),c,d}\left(1-\frac{y}{2N},1-\frac{x}{2N}\right)\frac{\pi_{q^{t}}^{(N),c}(1-\frac{y}{2N})}{\pi_{q^{s}}^{(N),d}(1-\frac{x}{2N})},$ $x\in\hat{S}_{s}^{(N),d},y\in\hat{S}_{t}^{(N),c},$
$\hat{\pi}_{s,t;v,w,u,r}^{(N),c,d}(x,y)$ $:=\pi_{q^{t},q^{s}}^{(N),d,c}\left(1-\frac{y}{2N},1-\frac{x}{2N}\right)\frac{\pi_{q^{t}}^{(N),d}(1-\frac{y}{2N})}{\pi_{q^{s}}^{(N),c}(1-\frac{x}{2N})},$ $x\in\hat{S}_{s}^{(N),c},y\in\hat{S}_{t}^{(N),d},$
$\hat{\pi}_{s,t}^{(N),d,d}(x,y)$ $:=\pi_{q^{t},q^{s}}^{(N),d,d}\left(1-\frac{y}{2N},1-\frac{x}{2N}\right)\frac{\pi_{q^{t}}^{(N),d}(1-\frac{y}{2N})}{\pi_{q^{s}}^{(N),d}(1-\frac{x}{2N})},$ $x\in\hat{S}_{s}^{(N),d},y\in\hat{S}_{t}^{(N),d}.$
The discrete support of $\pi_{q^{t},q^{s}}^{(N)}\left(1-\frac{y}{2N},\cdot\right)$ is not equal to the set $\hat{S}_{s}^{(N)}$, and likewise for $\pi_{q^{t},q^{s}}\left(\cdot,1-\frac{x}{2N}\right)$. These transition probabilities are set to be zero when $x$ and $y$ are not located, respectively, in $\hat{S}_{s}^{(N)}$ or $\hat{S}_{t}^{(N)}$.
The tangent process $\hat{\mathbb{Y}}_{s}^{(N)}$ has a symmetry on the continuous support, given by
(2.5) $\hat{\pi}_{t;v,w,u,r}^{(N),c}(y)=\hat{\pi}_{t;w,v,r,u}^{(N),c}(4N-y).$
The Askey-Wilson has a limiting process called the continuous dual Hahn process, which was defined by Corwin and Knizel [CK24] in order to characterize the stationary measure of the open KPZ.
2.3. The Continuous Dual Hahn Process
In this section, we give the definition of the continuous dual Hahn process. The definition in this paper is restricted to the time interval $t\in[0,C_{u,r})$ where $C_{u}=2r$ if $u\leq 0$ or $u\geq 1$ and $C_{u}=\min{\{2u,2r\}}$ if $u\in(0,1)$. It is possible to define the continuous dual Hahn process for all $t\in\mathbb{R}$, but the more general formulas are significantly more complicated, and not necessary for the arguments in this paper. In definitions which follow, $\mathfrak{p}_{t}(\cdot)$ will not be a probability distribution, but instead a positive distribution of infinite mass, while $\mathfrak{p}_{s,t}(\cdot,\cdot)$ will be a probability distribution.
Definition 2.2 (Section 6.2 [CK24]).
The continuous dual Hahn process is the process with one-point and transition measures given by
[…]
where $\mathfrak{p}^{c}_{t}(\cdot)$ is the one-point continuous dual Hahn density on the continuous support of the process, and $\mathfrak{p}^{d}_{t}(\cdot)$ is likewise the discrete mass defined on the discrete support of the process, and $\textsc{CDH}(V;a(x),b(x),c(x);\mathcal{F})$ is the continuous dual Hahn measure (See 2.3) for $x\in\mathcal{F}$ and $a(x),b(x),c(x)$ constants depending on $x$. These measures are supported on a set with a continuous component, $S^{c}_{t}$, and a discrete component, $S^{d}_{t}$, which is further broken up into two parts. We define these supports as
$S^{c}_{t}:=\mathbb{R}_{>0}$ $S^{d,u}_{t}:=\bigcup_{j=0}^{\lfloor-u+\frac{t}{2}\rfloor}\{x_{j}^{u}(t)\},$ $S_{t}^{d,v}:=\bigcup_{j=0}^{\lfloor-v-\frac{t}{2}\rfloor}\{x_{j}^{v}(t)\}.$
The first of the discrete collections is nonempty only when $u-t/2<0$, and the latter is nonempty only when $v+t/2<0$. The atoms in these sets are defined as
$x_{j}^{u}(t):=-4\left(u+j-\frac{t}{2}\right)^{2},$ $x_{j}^{v}(t):=-4\left(v+j+\frac{t}{2}\right)^{2}.$
On the continuous support, we define
(2.7) $\mathfrak{p}_{t}^{c}(x):=\frac{(u+v+1)(u+v)}{8\pi}\cdot\frac{|\Gamma\left(v+\frac{t}{2}+i\frac{\sqrt{x}}{2}\right)\Gamma\left(u-\frac{t}{2}+i\frac{\sqrt{x}}{2}\right)|^{2}}{\sqrt{x}|\Gamma(i\sqrt{x})|^{2}}\mathbf{1}_{x>0}.$
On the discrete support, we define
(2.8) $\mathfrak{p}_{t}^{d}(x_{j}^{u}(t)):=\frac{\Gamma(v-u+t)\Gamma(v+u+2)}{\Gamma(-2u+t)}\cdot\frac{\left(u+j-\frac{t}{2}\right)[2u-t,v+u]_{j}}{\left(u-\frac{t}{2}\right)[1,1-v+u-t]_{j}},$
(2.9) $\mathfrak{p}_{t}^{d}(x_{j}^{v}(t)):=\frac{\Gamma(u-v-t)\Gamma(v+u+2)}{\Gamma(-2v-t)}\cdot\frac{\left(v+j+\frac{t}{2}\right)[2v+t,v+u]_{j}}{\left(v+\frac{t}{2}\right)[1,1-v+u-t]_{j}}.$
To define the transition densities, we use $\mathcal{S}_{t}$ to denote the union of the continuous and discrete supports at time $t$. In the definition which follows, we use the notation $\textsc{CDH}(V;a,b,c;\mathcal{F})$ to denote the continuous dual Hahn measure, which is introduced in 2.3. For $x\in S_{s}^{c}$, we define
(2.10) $\mathfrak{p}_{s,t}(x,V):=\textsc{CDH}\left(V;u-\frac{t}{2},\frac{t-s}{2}+i\frac{\sqrt{x}}{2},\frac{t-s}{2}-i\frac{\sqrt{x}}{2};\mathcal{S}_{t}\right).$
If $u-\frac{s}{2}<0$ (and consequently, $S_{s}^{d}=S_{s}^{d,u}$, then for $j\in[[0,\lfloor-u+\frac{s}{2}\rfloor]]$,
(2.11) $\mathfrak{p}_{s,t}(x_{j}^{u}(s),V):=\textsc{CDH}\left(V;u-\frac{t}{2},-u+\frac{t}{2}-j,u+\frac{t}{2}-s+j;\mathcal{S}_{t}\right).$
If $v+\frac{s}{2}<0$ (and consequently, $S^{d}_{s}=S_{s}^{d,v}$, then for $j\in[[0,\lfloor-v-\frac{s}{2}\rfloor]]$,
(2.12) $\mathfrak{p}_{s,t}(x_{j}^{v}(s),V):=\textsc{CDH}\left(V;v+j+\frac{t}{2},\frac{t}{2}-s-v-j,u-\frac{t}{2};\mathcal{S}_{t}\right).$
In all other cases, we define $\mathfrak{p}_{s,t}$ to be zero.
Definition 2.3 (Definitions 6.3, 6.4 [CK24]).
Assume $a,b,c\in\mathbb{R}$ with $a+b,a+c>0$ or $a\in\mathbb{R}$ and $b=\overline{c}$ with $Re(b)=Re(c)>0$. For such $a,b,c$ we define a density function
$\textsc{CDH}^{c}(x;a,b,c):=\frac{1}{8\pi}\cdot\frac{\bigg{|}\Gamma\left(a+i\frac{\sqrt{x}}{2}\right)\Gamma\left(b+i\frac{\sqrt{x}}{2}\right)\Gamma\left(c+i\frac{\sqrt{x}}{2}\right)\bigg{|}^{2}}{\Gamma(a+b)\Gamma(a+c)\Gamma(b+c)\sqrt{x}|\Gamma(i\sqrt{x})|^{2}}\mathbf{1}_{x>0}.$
Similarly, in the setting where $a<0$ and $\mathcal{F}$ is a finite subset of $\mathbb{R}$ with size $|\mathcal{F}|=\lfloor-a\rfloor+1$ and elements $x_{0}<x_{1}<\cdot\cdot\cdot<x_{\lfloor-a\rfloor}$, then assume that $b,c,$ and $j$ satisfy one of the following conditions:
  1. (1)
    $b=\overline{c}$ with $Re(b),Re(c)>0$ and $j\in[[0,\lfloor-a\rfloor]]$
  2. (2)
    $b,c\in\mathbb{R}$ with $a+b,a+c>0$ and $j\in[[0,\lfloor-a\rfloor]]$
  3. (3)
    $b,c\in\mathbb{R}$ and $b,b+c,c-a>0$ with $a+b=-k$ for $k\in\mathbb{Z}_{\geq 0}$ and $j\in[[0,k]]$.
When one of these conditions is satisfied, we define the discrete mass at $x_{j}$ to be
$\textsc{CDH}^{d}(x_{j};a,b,c;\mathcal{F}):=\frac{[2a,a+b,a+c]_{j}(a+j)\Gamma(b-a)\Gamma(c-a)}{[1,a-b+1,a-c+1]_{j}a\Gamma(-2a)\Gamma(b+c)}(-1)^{j}.$
On all other arguments, $\textsc{CDH}^{d}(\cdot;a,b,c;\mathcal{F})$ is zero. Finally, we define the full continuous dual Hahn measure $\textsc{CDH}(V;a,b,c;\mathcal{F})$ in the following three cases:
  • For $a\geq 0$; $\mathcal{F}=\emptyset$ and either $b=\overline{c}$ with $Re(b),Re(c)>0$ or $b,c\in\mathbb{R}_{>0}$,
[…]
  • For $a<0$; $\mathcal{F}$ a finite subset of $\mathbb{R}$ with size $|\mathcal{F}|=\lfloor-a\rfloor+1$ and elements $x_{0}<x_{1}<\cdot\cdot\cdot<x_{\lfloor-a\rfloor}$, and either $b=\overline{c}$ with $Re(b),Re(c)>0$ or $b,c\in\mathbb{R}$ with $a+b,a+c>0$,
[…]
  • For $a<0$; $\mathcal{F}$ a finite subset of $\mathbb{R}$ with size $|\mathcal{F}|=\lfloor-a\rfloor+1$ and elements $x_{0}<x_{1}<\cdot\cdot\cdot<x_{\lfloor-a\rfloor}$ and $b,c\in\mathbb{R}$ with $b,b+c,c-a>0$ and $a+b=-k$ for $k\in\mathbb{Z}_{\geq 0}$
[…]
Finally, we note the following property of the continuous dual Hahn process.
Lemma 2.4 (Lemma 6.9 [CK24]).
For $u,v\in\mathbb{R}$ with $u+v>0$ and $0\leq s<t<w<C_{u,v}$ and for any Borel set $V\subset\mathbb{R}$,
$\int_{\mathbb{R}}\mathfrak{p}_{s}(dm)\mathfrak{p}_{s,t}(m,V)=\mathfrak{p}_{t}(V)$ $\int_{\mathbb{R}}\mathfrak{p}_{s,t}(m,dr)\mathfrak{p}_{t,w}(r,V)=\mathfrak{p}_{s,w}(m,V)$
2.4. Notation for Asymptotics
We will use the following notation to simplify the expressions for the logarithms of the $q$-pochammer symbols.
(2.13) $\mathcal{A}^{+}[\kappa,z]=-\frac{\pi^{2}}{6\kappa}-\left(z-\frac{1}{2}\right)\log{(\kappa)}-\log{\left[\frac{\Gamma(z)}{\sqrt{2\pi}}\right]}$ $\mathcal{A}^{-}[\kappa,z]=\frac{\pi^{2}}{12\kappa}-\left(z-\frac{1}{2}\right)\log{(2)}.$
The following lemma will also be crucial to our asymptotic analysis.
Lemma 2.5 (Proposition 2.3 [CK24]).
For $\kappa\in(0,1)$, $z\in\mathbb{C}$, and $q=\exp{(-\kappa)}$, and $m\in\mathbb{Z}_{>0}$,
$\log{\left[(q^{z};q)_{\infty}\right]}=\mathcal{A}^{+}[\kappa,z]-\sum_{n=1}^{m-1}\frac{B_{n+1}(z)B_{n}}{n(n+1)!}\kappa^{n}+\textsc{Error}^{+}_{m}[\kappa,z],$
$\log{\left[(-q^{z};q)_{\infty}\right]}=\mathcal{A}^{-}[\kappa,z]-\sum_{n=1}^{m-1}(2^{n}-1)\frac{B_{n+1}(z)B_{n}}{n(n+1)!}\kappa^{n}+\textsc{Error}^{-}_{m}[\kappa,z],$
where $B_{n}(z)$ and $B_{n}$ denote Bernoulli polynomials and Bernoulli numbers, respectively, and $\textsc{Error}_{m}^{\pm}[\kappa,z]$ are error terms. For any $\alpha\in(0,\pi)$, $\epsilon\in(0,1/2)$ and any $b\in(m-1,m)$ there exist $C,\kappa_{0}>0$ such that for all $\kappa\in(0,\kappa_{0})$ and all $z\in\mathbb{C}$ with $|Im(z)|<\frac{\alpha}{\kappa}$,
(2.14) $|\textsc{Error}_{1}^{\pm}[\kappa,z]|\leq C(\kappa(1+|z|)^{2}+\kappa^{b}(1+|z|)^{1+2b+\epsilon}).$
[…]
We also use the following notation as shorthand
$\mathcal{A}^{\pm}_{N}[z_{1},...,z_{k}]=\sum_{i=1}^{k}\mathcal{A}^{\pm}\left[\kappa,z_{i}\right],$ $E^{\pm}_{N}[z_{1},...,z_{k}]=\sum_{i=1}^{k}\textsc{Error}_{1}^{\pm}\left[\kappa,z_{i}\right].$
2.5. Convergence of the Laplace transform
Finally, we state the result about pointwise convergence of Laplace transforms, which we will use to obtain our weak convergence.
Proposition 2.6 (Section 5.14, page 378, (5.14.8) [HJ94]).
For a collection of random vectors of length $d$, $\{\mathbf{X}^{(i)}\}_{i\in\mathbb{N}}$, with $\mathbf{X}^{(i)}:=(X^{(i)}_{1},\ldots,X^{(i)}_{d})$ and $\mathbf{z}\in\mathbb{R}^{d}$ defined so that $\mathbf{z}:=(z_{1},\ldots,z_{d})$, we define the collection of functions $\{L_{i}(\mathbf{z})\}_{i\in\mathbb{N}}$ to be the Laplace transforms of each of the $\mathbf{X}^{(i)}$, respectively,
$L_{i}(\textbf{z})=\mathbb{E}\left[\exp{\left(\sum_{j=1}^{d}z_{j}X_{j}^{(i)}\right)}\right].$
[…]
3. Convergence of the Askey-Wilson Densities
The purpose of this section and the next is to demonstrate that the sequence of Laplace transforms of the Askey-Wilson process, under the scaling assumptions without Liggett’s condition, given in 1.1, converges pointwise to the Laplace transform of the open KPZ stationary measure.
Proposition 3.1.
[…]
(3.1) $\varphi^{(N)}_{u,v,w,r}(\textbf{c},\textbf{x})$ $=\frac{\mathbb{E}\left[\prod_{k=1}^{d+1}\left(\cosh(s_{k}^{(N)})+\mathbb{Y}_{e^{-2s_{k}^{(N)}}}\right)^{n_{k}-n_{k-1}}\right]}{\mathbb{E}\left[(1+\mathbb{Y}_{1})^{N}\right]}.$
For all $\textbf{c}\in(0,C_{u,r})^{d}$, defined so that
$C_{u,r}$ :=$:=\begin{cases}2&u\leq 0,r\leq 0\\ 2r&r>0,u\leq 0\\ \min{\{2r,2u\}}&u,r>0\end{cases}$
, and $u+v>0$ and $w,r>0$, this sequence converges point-wise to a function $\varphi_{u,v}(\textbf{c},\textbf{x})$, defined
(3.2) $\varphi_{u,v}(\textbf{c},\textbf{x})$ $:=\frac{\mathbb{E}\left[\exp{\left(\frac{1}{4}\sum_{k=1}^{d+1}(s_{k}^{2}-\mathbb{T}_{s_{k}})(x_{k}-x_{k-1})\right)}\right]}{\mathbb{E}\left[\exp{\left(-\frac{1}{4}\mathbb{T}_{0}\right)}\right]},$
where $\mathbb{T}_{t}$ is the continuous dual Hahn process. Moreover, $\varphi_{u,v}(\textbf{c},\textbf{x})$ is the Laplace transform of the stationary measure of the open KPZ equation constructed in [CK24].
The first step in the proof of 3.1 is to demonstrate the convergence of the Askey-Wilson process measures to those of the continuous dual Hahn process, which we do in the rest of this section. In Section 4 we will complete the proof of 3.1. Finally, in Section 5, we give the proof of 1.3.
We will use the notation $x_{j}^{\chi}(t)$ and $y_{j}^{(N),\chi}$, respectively, for the locations of the atoms of the continuous dual Hahn process and Askey-Wilson process (at the $N$-th step of scaling) associated to $\chi\in\{q^{v+t/2},q^{u-t/2},-q^{w+t/2},-q^{r-t/2}\}$.
Lemma 3.2.
Under the scaling of the Askey-Wilson measure when $q=\exp{\left(\frac{-2}{\sqrt{N}}\right)}$, see 2.1,
  1. (1)
    For all $y\in\hat{S}^{c}$ and $t\in\mathbb{R}_{>0}$, for all $\eta>1$ there exist $N_{0}\in\mathbb{N}$ and $C,\chi>0$ so that for all $N>N_{0}$ and $y\in[0,4N]$,
    $N^{u+v}\sqrt{1-\frac{y}{4N}}\hat{\pi}_{t}^{(N),c}(y)=C_{u,v,w,r}\cdot\mathfrak{p}_{t}^{c}(y)\cdot e^{\textsc{Error}_{t}^{(N),c}(y)},$
    with
    $|\textsc{Error}_{t}^{(N),c}(y)|\leq CN^{-\chi}(1+\sqrt{r})^{\eta}.$
    In particular, we have
    $\lim_{N\to\infty}N^{u+v}\hat{\pi}_{t}^{(N),c}(y)$ $=C_{u,v,w,r}\cdot\mathfrak{p}_{t}^{c}(y),$
    where
    $C_{u,v,w,r}:=\frac{\Gamma(u+v+w+r)}{\Gamma(u+v+2)\Gamma(w+r)}.$
  2. (2)
    Similarly, for $\hat{y}_{j}^{(N),\chi}(t)\in\hat{S}^{d}_{t}$ and $t\in(-2w,2r)$, there exist $C_{1},C_{2},\eta_{1},\eta_{2}\in\mathbb{R}_{>0}$ such that for all $N\in\mathbb{Z}_{\geq 1}$, the atoms $y_{j}^{(N),\chi}$ for $\chi\in\{q^{v+\frac{t}{2}},q^{u-\frac{t}{2}}\}$ satisfy
    $N^{u+v}\hat{\pi}_{t}^{(N),d}(\hat{y}^{(N),v}(t))=C_{u,v,w,r}\cdot\mathfrak{p}_{t}^{d}(x_{j}^{v}(t))\cdot e^{\textsc{Error}_{t}^{(N),d}(x_{j}^{v}(t))},$
    $N^{u+v}\hat{\pi}_{t}^{(N),d}(\hat{y}^{(N),u}(t))=C_{u,v,w,r}\cdot\mathfrak{p}_{t}^{d}(x_{j}^{u}(t))\cdot e^{\textsc{Error}_{t}^{(N),d}(x_{j}^{u}(t))},$
    with the exponential error terms bounded by
    $|\textsc{Error}_{t}^{(N),d}(x_{j}^{v}(t))|\leq C_{1}N^{-\eta_{1}},$
    $|\textsc{Error}_{t}^{(N),d}(x_{j}^{u}(t))|\leq C_{2}N^{-\eta_{2}},$
    and $C_{u,v,w,r}$ as defined above. In particular, we have
    $\lim_{N\to\infty}N^{u+v}\hat{\pi}_{t}^{(N),d}(\hat{y}_{j}^{(N),\chi}(t))=C_{u,v,w,r}\cdot\mathfrak{p}_{t}^{d}(x_{j}^{\chi}(t)).$
The limiting measure $\mathfrak{p}_{t}^{c}(y)$ and $\mathfrak{p}_{t}^{d}(y)$ are the one-point measures of the continuous dual Hahn process, see Section 2.3.
Similarly, we can compute the limit of $\hat{\pi}_{s,t}^{(N),c,c}(\cdot,\cdot)$ and $\hat{\pi}_{s,t}^{(N),d,c}(\cdot,\cdot)$ under scaling.
Lemma 3.3.
With scaling conditions as in 1.1, and for all $s<t$
  1. (1)
    For all $m,r\in\hat{S}^{(N),c}=[0,4N]$ and $\eta>0$, there exists $N_{0}\in\mathbb{Z}_{\geq 1}$ and $C,\chi\in\mathbb{R}_{>0}$ such that for all $N>N_{0}$
    $\sqrt{1-\frac{r}{4N}}\hat{\pi}_{s,t}^{(N),c,c}(m,r)=\mathfrak{p}_{s,t}^{c,c}(m,r)\cdot e^{\textsc{Error}_{s,t}^{(N),c,c}(m,r)}$
    with an error term satisfying
    $|\textsc{Error}_{s,t}^{(N),c,c}(m,r)|\leq CN^{-\chi}(1+\sqrt{m})^{\eta}+CN^{-\chi}(1+\sqrt{r})^{\eta}.$
    In particular,
    $\lim_{N\to\infty}\hat{\pi}_{s,t}^{(N),c,c}(m,r)=\mathfrak{p}_{s,t}^{c,c}(m,r).$
  2. (2)
    If $s,t\in(-2b,2d)$ and $v+\frac{s}{2}<0$ so that $\hat{S}_{s}^{(N),d}$ is entirely composed of $v$ atoms $\hat{y}_{j}^{(N),v}(s)$ for $j\in[[0,\lfloor-v-\frac{s}{2}\rfloor]]$ then for each such $j$, and all $r\in\hat{S}^{(N),c}=[0,4N]$ there exists $N_{0}\in\mathbb{Z}_{\geq 1}$ and $C,\chi\in\mathbb{R}_{>0}$ such that for all $N>N_{0}$
    $\hat{\pi}_{s,t}^{(N),d,c}\left(\hat{y}_{j}^{(N),v}(s),r\right)=\mathfrak{p}_{s,t}^{d,c}(x_{j}^{v}(s),r)\cdot e^{\textsc{Error}_{s,t}^{(N),d,c}(x_{j}^{v}(s),r)}$
    with
    $|\textsc{Error}_{s,t}^{(N),d,c}(x_{j}^{v}(s),r)|\leq CN^{-\chi}(1+\sqrt{r})^{\eta}.$
    In particular,
    $\lim_{N\to\infty}\hat{\pi}_{s,t}^{(N),d,c}(\hat{y}_{j}^{(N),v}(s),r)=\mathfrak{p}_{s,t}^{d,c}(x_{k}^{v}(s),r).$
  3. (3)
    If $s,t\in(-2w,2r)$ and $u-\frac{s}{2}<0$ so that $\hat{S}_{s}^{(N),d}$ is entirely composed of $u$ atoms $\hat{y}_{j}^{(N),u}(s)$ for $j\in[[0,\lfloor-u+\frac{s}{2}\rfloor]]$ then for each such $j$ and for all $r\in\hat{S}^{(N),c}=[0,4N]$,
    $\lim_{N\to\infty}\hat{\pi}_{s,t}^{(N),d,c}(\hat{y}_{j}^{(N),v}(s),r)=0$
Finally, we obtain the following asymptotic result for $\hat{\pi}_{t}^{(N),d,d}(\cdot,\cdot)$ and $\hat{\pi}_{t}^{(N),c,d}(\cdot,\cdot)$.
Lemma 3.4.
With scaling conditions as defined in 1.1, and for all $s,t\in[0,C_{u,r})$ with $s<t$,
  1. (1)
    If $v+\frac{s}{2}<0$ then the discrete support of $\hat{\pi}_{s,t}^{(N),d,d}(\hat{y}_{j}^{(N),v}(s),\cdot)$ coincides with the points $\hat{y}_{k}^{(N),v}(t)$ for $k\in[[0,\lfloor-v-\frac{t}{2}\rfloor]]$ and $\hat{y}_{k}^{(N),u}(t)$ for $k\in[[0,\lfloor-u+\frac{t}{2}\rfloor]]$ and there exist $C,\chi>0$ such that for all $N>0$ and $k\in[[0,\lfloor-v-\frac{t}{2}\rfloor]],$ respectively, $k\in[[0,\lfloor-u+\frac{t}{2}\rfloor]]$
    $\hat{\pi}_{s,t}^{(N),d,d}(\hat{y}_{j}^{(N),v}(s),\hat{y}_{k}^{(N),v}(t))=\mathfrak{p}_{s,t}^{d,d}(x_{j}^{v}(s),x_{k}^{v}(t))\cdot e^{\textsc{Error}_{s,t}^{(N),d,d}(v,v)}$
    $\hat{\pi}_{s,t}^{(N),d,d}(\hat{y}_{j}^{(N),v}(s),\hat{y}_{k}^{(N),u}(t))=\mathfrak{p}_{s,t}^{d,d}(x_{j}^{v}(s),x_{k}^{u}(t))\cdot e^{\textsc{Error}_{s,t}^{(N),d,d}(v,u)}$
    and the error terms are bounded so that
    $|\textsc{Error}_{s,t}^{(N),d,d}(\cdot,\cdot)|\leq CN^{-\chi}.$
    In particular,
    $\lim_{N\to\infty}\hat{\pi}_{s,t}^{(N),d,d}(\hat{y}_{j}^{(N),v}(s),\hat{y}_{k}^{(N),v}(t))=\mathfrak{p}_{s,t}^{d,d}(x_{j}^{v}(s),x_{k}^{v}(t)).$
  2. (2)
    If $u-\frac{s}{2}<0$ then the discrete support of $\hat{\pi}_{s,t}^{(N),d,d}(\hat{y}_{j}^{(N),u}(s),\cdot)$ does not contain any atoms.
  3. (3)
    If $u-\frac{s}{2}<0$ then the discrete support of $\hat{\pi}_{s,t}^{(N),d,d}(\hat{y}_{j}^{(N),u}(s),\cdot)$ coincides with the points $\hat{y}_{k}^{(N),v}(t)$ for $k\in[[0,\lfloor-v-\frac{t}{2}\rfloor]]$ and
    $\lim_{N\to\infty}\hat{\pi}_{s,t}^{(N),d,d}(\hat{y}_{j}^{(N),u}(s),\hat{y}_{k}^{(N),v}(t))=\mathfrak{p}_{s,t}^{d,d}(x_{j}^{u}(s),x_{k}^{v}(t))$
  4. (4)
    If $v+\frac{s}{2}<0$ and $r\in\hat{S}^{(N),c}=[0,4N]$ then the Askey-Wilson measure $\hat{\pi}_{s,t}^{(N),c,d}(r,\cdot)$ has no atomic part. When $u-\frac{s}{2}<0$ and $r\in\hat{S}^{(N),c}=[0,4N]$, the measure $\hat{\pi}_{s,t}^{(N),c,d}(r,\cdot)$ has discrete support on the atoms $\hat{y}_{k}^{(N),u}(t)$ and there exist $C,\chi\in\mathbb{R}_{>0}$ such that for all $N\in\mathbb{Z}_{\geq 1}$,
    $\hat{\pi}_{s,t}^{(N),c,d}\left(r,\hat{y}_{k}^{(N),u}(t)\right)=\mathfrak{p}_{s,t}^{c,d}(r,x_{k}^{u}(t))\cdot e^{\textsc{Error}_{s,t}^{c,d}(r,x_{k}^{u}(t))}$
    with the error bounded
    $|\textsc{Error}_{s,t}^{c,d}(r,x_{k}^{u}(t))|\leq CN^{-\chi}.$
    In particular,
    $\lim_{N\to\infty}\hat{\pi}_{s,t}^{(N),c,d}(r,\hat{y}_{k}^{(N),u}(t))=\mathfrak{p}_{s,t}^{c,d}(r,x_{k}^{u}(t)).$
3.1. Convergence of One-point Distributions
This section deals with the proof of Lemma 3.2.
Proof of Lemma 3.2 (1).
We begin by reminding ourselves of the expression for $\hat{\pi}_{t;v,w,u,r}^{(N),c}(\cdot)$. As before, we set $q=\exp{\left(-\frac{2}{\sqrt{N}}\right)}$. We will deal with the expression for the rescaled, time-reversed process $\hat{\mathbb{Y}}_{t}$, and the corresponding one-point measures $\hat{\pi}_{t;v,w,u,r}^{(N),c}(\cdot)$.
(3.3) $\hat{\pi}_{t;v,w,u,r}^{(N),c}(y)$ $=\frac{1}{2N}\pi_{q^{t}}^{(N),c}\left(1-\frac{y}{2N}\right)$
$=\frac{1}{4N\pi(q^{u+v+w+r})_{\infty}\sqrt{\frac{y}{N}}\sqrt{1-\frac{y}{4N}}}$
×$\times\frac{(q,-q^{v+w+t},q^{u+v},-q^{v+r},-q^{u+w},q^{w+r},-q^{u+r-t})_{\infty}|(q^{i\sqrt{N}\theta})_{\infty}|^{2}}{|(q^{v+\frac{t}{2}+i\frac{\sqrt{N}\theta}{2}},-q^{w+\frac{t}{2}+i\frac{\sqrt{N}\theta}{2}},q^{u-\frac{t}{2}+i\frac{\sqrt{N}\theta}{2}},-q^{r-\frac{t}{2}+i\frac{\sqrt{N}\theta}{2}})_{\infty}|^{2}}$
We will only prove a bound in the interval $y\in[0,2N]$. The way that this bound can be applied to prove the full Lemma 3.2 is by using the symmetry (2.5) to turn this bound into in the proof by first proving our desired bound for $\hat{\pi}^{(N),c}_{t;v,w,u,r}(y)$ on the interval $y\in[0,2N]$, and then noting that the desired bound on the interval $[2N,4N]$ is significantly weaker than the bound on the interval $[0,2N]$ which arises from $\hat{\pi}^{(N),c}_{t;w,v,r,u}(4N-y)$. Therefore, we first prove the appropriate bound on (3.3) in the interval $r\in[0,2N]$, and then use this identity to reach the remaining $r\in[2N,4N].$ We take the logarithm of the one-point measure,
(3.4) $\log{\left(\sqrt{1-\frac{y}{4N}}\hat{\pi}_{t;v,w,u,r}^{(N),c}(y)\right)}$ $=-\log{\left(4\pi N\sqrt{\frac{y}{N}}\right)}+\mathcal{A}(\sqrt{y})+E^{\mathcal{A}}_{N}(y)+E^{\mathcal{E}}_{N}(y)$
where, in the expression above,
$\mathcal{A}_{N}(\sqrt{y})$ $:=\mathcal{A}^{+}_{N}\left[1,u+v,w+r,i\frac{\sqrt{y}}{2},-i\frac{\sqrt{y}}{2}\right]$
$+\mathcal{A}^{-}_{N}\left[v+w+t,v+r,u+w,u+r-t\right]$
$-\mathcal{A}_{N}^{+}\left[u+v+w+r,v+i\frac{\sqrt{y}}{2}+\frac{t}{2},u+i\frac{\sqrt{y}}{2}-\frac{t}{2},v-i\frac{\sqrt{y}}{2}+\frac{t}{2},u-i\frac{\sqrt{y}}{2}-\frac{t}{2}\right]$
$-\mathcal{A}_{N}^{-}\left[w+i\frac{\sqrt{y}}{2}+\frac{t}{2},r+i\frac{\sqrt{y}}{2}-\frac{t}{2},w-i\frac{\sqrt{y}}{2}+\frac{t}{2},r-i\frac{\sqrt{y}}{2}-\frac{t}{2}\right].$
$E^{\mathcal{A}}_{N}(y)$ $:=\mathcal{A}_{N}(\sqrt{y}+E^{\theta}_{N}(y))-\mathcal{A}_{N}(\sqrt{y}),$
$E_{N}^{\mathcal{E}}(y)$ $:=\mathcal{E}_{N}(\sqrt{y}+E_{N}^{\theta}(y)),$
$E_{N}^{\theta}(y)$ $:=\sqrt{N}\arccos{\left(1-\frac{y}{2N}\right)}-\sqrt{y}.$
are a concise way of containing the logarithm of the q-pochammer symbols in $\hat{\pi}_{t;v,w,u,r}^{(N),c}(y)$. The notation $\mathcal{A}^{+}$ and $\mathcal{A}^{-}$ was defined in (2.13), and $E_{N}^{\theta}(y)$ arises as a concise expression for the error terms in the Taylor expansion $\sqrt{N}\theta_{y}^{(N)}=\sqrt{y}+E_{N}^{\theta}(y)$.
We can simplify the first two terms on the right-hand side of (3.4) considerably by expanding $\mathcal{A}(\sqrt{y})$ through the definitions of $\mathcal{A}^{+}[z_{1},...,z_{k}]$ and $\mathcal{A}^{-}[z_{1},...,z_{k}]$ from (2.13), and then canceling where appropriate. The result of that calculation yields
$-\log{\left(4\pi N\sqrt{\frac{y}{N}}\right)}+\mathcal{A}_{N}(\sqrt{y})$ $=-\left(u+v\right)\log{(N)}$
$+\log{\left(\frac{\Gamma\left(u+v+w+r\right)\bigg{|}\Gamma\left(v+\frac{i\sqrt{y}}{2}+\frac{t}{2}\right)\bigg{|}^{2}\bigg{|}\Gamma\left(u+\frac{i\sqrt{y}}{2}-\frac{t}{2}\right)\bigg{|}^{2}}{8\pi\sqrt{y}\Gamma(u+v)\Gamma\left(w+r\right)|\Gamma(i\sqrt{y}/2)|^{2}}\right)}.$
Using the definitions of $\mathfrak{p}_{t}^{c}(y)$ and $C_{u,v,w,r}$ from Lemma 3.2, we can rewrite (3.4) as
$\log\left(N^{u+v}\sqrt{1-\frac{y}{4N}}\hat{\pi}_{t;v,w,u,r}^{(N),c}(y)\right)$ $=\log{(\mathfrak{p}_{t}^{c}(y))}+\log{(C_{u,v,w,r})}+E_{N}^{\mathcal{A}}(y)+E_{N}^{\mathcal{E}}(y).$
What remains is to check that the error terms are bounded appropriately. These arguments are identical to those in [CK24], except that the error terms may now vary in $w$ and $r$. We summarize them here for completeness. First we will quote, as a lemma, a few intermediate results from [CK24], which provide helpful bounds.
Lemma 3.5 (p. 53-54, [CK24]).
  1. (1)
    For all $\eta>1$ there exists $N_{0}\in\mathbb{Z}_{\geq 1}$ and $C,\chi>0$ such that for all $N>N_{0}$ and $y\in\hat{S}^{(N),c}=[0,2N]$,
    $E_{N}^{\theta}(y)\log{(N)}\leq CN^{-\chi}(1+\sqrt{y})^{\eta}.$
  2. (2)
    For any fixed $z$ and for all $b\in(0,1)$ there exist $\kappa_{0}>0$ and $C>0$ such that for all $\kappa<\kappa_{0}$,
    $|\textsc{Error}_{1}^{\pm}[\kappa,z]|\leq C(\kappa+\kappa^{b})\leq C^{\prime}N^{-\frac{b}{2}}.$
  3. (3)
    For $z_{c}(y)=c\pm i\frac{\sqrt{y}+E_{N}^{\theta}(y)}{2}$, for any $b\in(0,1)$ and $\epsilon\in(0,1/2)$, there exists $C>0$ and $N_{0}\in\mathbb{Z}_{\geq 1}$ such that for all $N>N_{0}$ and $y\in\hat{S}^{(N),c}=[0,2N]$,
    $|\textsc{Error}_{1}^{\pm}[\kappa,z_{c}(y)]|\leq C\left(N^{-1/2}(1+\sqrt{y})^{2}+N^{-b/2}(1+\sqrt{y})^{1+2b+\epsilon}\right).$
  4. (4)
    For $z_{c}(y)=c\pm i\frac{\sqrt{y}+E_{N}^{\theta}(y)}{2}$, $z^{\prime}_{c}(y)=c\pm i\frac{\sqrt{y}}{2}$ and for any $c\in\mathbb{R}$ fixed, there exists $N_{0}\in\mathbb{Z}_{\geq 1}$ and $C,\chi>0$ such that for all $N>N_{0}$ and $y\in[0,4N]$,
    $|\log{[\Gamma(z^{\prime}_{c}(y))]}-\log{[\Gamma(z_{c}(y))]}|\leq CN^{-\chi}(1+\sqrt{y})^{\eta}.$
With these results in hand, we can state the bound on the error terms.
Lemma 3.6.
For all $\eta>1$ there exists $N_{0}$ and $C,\chi>0$ such that for all $N>N_{0}$ and $y\in[0,2N]$,
(3.5) $|E_{N}^{\mathcal{A}}(y)|,|E_{N}^{\mathcal{E}}(y)|\leq CN^{-\chi}(1+\sqrt{y})^{\eta}$
Proof.
There are two types of terms involved in $E^{\mathcal{E}}_{N}(y)$. We separate the terms of $E_{N}^{\mathcal{E}}(y)$ into terms which only depend on $u,v$ and $t$ (type (1)) and terms which depend on $y$ (type (2)).
Type (1) $=E_{N}^{+}\left[1,u+v,w+r\right]+E_{N}^{-}\left[v+t+w,v+r,u+w,u+r-t\right]-E_{N}^{+}\left[u+v+w+r\right]$
Type (2) $=E_{N}^{+}\left[\frac{i\sqrt{N}}{2}\arccos{\left(1-\frac{y}{4N}\right)},\frac{-i\sqrt{N}}{2}\arccos{\left(1-\frac{y}{4N}\right)}\right]$
$-E_{N}^{+}\left[v+\frac{t}{2}+\frac{i\sqrt{N}}{2}\arccos{\left(1-\frac{y}{4N}\right)},v+\frac{t}{2}-\frac{i\sqrt{N}}{2}\arccos{\left(1-\frac{y}{4N}\right)}\right]$
$-E_{N}^{+}\left[u-\frac{t}{2}+\frac{i\sqrt{N}}{2}\arccos{\left(1-\frac{y}{4N}\right)},u-\frac{t}{2}-\frac{i\sqrt{N}}{2}\arccos{\left(1-\frac{y}{4N}\right)}\right]$
$-E_{N}^{-}\left[w+\frac{t}{2}+\frac{i\sqrt{N}}{2}\arccos{\left(1-\frac{y}{4N}\right)},w+\frac{t}{2}-\frac{i\sqrt{N}}{2}\arccos{\left(1-\frac{y}{4N}\right)}\right]$
$-E_{N}^{-}\left[r-\frac{t}{2}+\frac{i\sqrt{N}}{2}\arccos{\left(1-\frac{y}{4N}\right)},r-\frac{t}{2}-\frac{i\sqrt{N}}{2}\arccos{\left(1-\frac{y}{4N}\right)}\right]$
We note that the terms of type (1) do not vary in $y$ and we can therefore apply the bound from Lemma 3.5 (2). The terms of type (2) come from expressions of the form $\textsc{Error}_{1}^{\pm}[\kappa,z]$ with $z$ taking the form $z_{c}(y)=c+i\left(\frac{\sqrt{y}+E_{N}^{\theta}(y)}{2}\right)$ for $c\in\mathbb{R}$. Therefore, we can directly apply Lemma 3.5 (3) to bound these terms.
Turning our attention to $E^{\mathcal{A}}_{N}(y)$ (recall $E^{\mathcal{A}}_{N}(y)=\mathcal{A}_{N}(\sqrt{y}+E_{N}^{\theta}(y))-\mathcal{A}_{N}(\sqrt{y})$), and using the expressions for $\mathcal{A}^{\pm}[z_{1},..,z_{k}]$ from (2.13), we see that the subtraction in the definition of $E_{N}^{\mathcal{A}}$ causes all the terms involving a factor of $\pi^{2}$ to cancel. The terms which remain take one of three forms:
  1. (1)
    Terms associated to the $\left(z-\frac{1}{2}\right)\log{(\kappa)}$ term in $\mathcal{A}^{+}$:
    The part of these terms which remains after the cancellations takes the form $\frac{iE_{N}^{\theta}(y)}{2}\log{(\kappa)}$ which is proportional to $E_{N}^{\theta}(y)\log{(N)},$ and therefore, we can apply Lemma 3.5 to bound this contribution by $CN^{-\chi}(1+\sqrt{r})^{\eta}$ for a small enough $\chi$.
  2. (2)
    Terms associated to the $\log{\left[\frac{\Gamma(z)}{\sqrt{2\pi}}\right]}$ term in $\mathcal{A}^{+}:$
    The contributions from terms of this kind after cancellation take the form $\log{(\Gamma(z^{\prime}_{c}(y)))}-\log{(\Gamma(z_{c}(y)))}$ with
[…]
    Therefore, we can apply Lemma 3.5 (4) to get the appropriate bound.
  3. (3)
    Terms associated to the $\left(z-\frac{1}{2}\right)\log{(2)}$ term in $\mathcal{A}^{-}:$
    We note that whenever $N\geq 2$, $\log{(2)}\leq\log{(N)}$ and consequently, since terms of type (3) take the form $E_{N}^{\theta}(y)\log{(2)}$, we can apply Lemma 3.5 (1) to obtain the desired bound.
Therefore, using the symmetry of the tangent process (2.5), we conclude that for all $\eta>1$ there exist $N_{0}\in\mathbb{N}$ and $C,\chi>0$ so that for all $N>N_{0}$ and $y\in[0,4N]$,
(3.6) $N^{u+v}\sqrt{1-\frac{y}{4N}}\hat{\pi}_{t;v,w,u,r}^{(N),c}(y)=\mathfrak{p}_{t}^{c}(y)\cdot e^{\textsc{Error}_{t}^{(N),c}(y)},$
with $|\textsc{Error}_{t}^{(N),c}(y)|\leq CN^{-\chi}(1+\sqrt{r})^{\eta}.$ This allows us to complete the proof of Lemma 3.2 (1), since
$\lim_{N\to\infty}N^{u+v}\hat{\pi}_{t;w,v,u,r}^{(N),c}(y)=C_{b,d,u,v}\cdot\mathfrak{p}_{t}^{c}(y).$
We continue to prove Lemma 3.2 (2). In order to prove the lemma, we need to show convergence of the atoms’ locations and weights. The convergence result for the locations of the atoms is unchanged from that stated in [CK24], as the locations of the atoms associated to $\chi\in\{q^{v+\frac{t}{2}},q^{u-\frac{t}{2}}\}$ do not depend on parameters $w$ and $r$, and our choice of time interval excludes all other atoms.
Lemma 3.7 (Lemma 9.3 [CK24]).
Assume $t\in(-2w,2r)$. The location of the atoms of $\hat{\pi}_{t}^{(N),d}(\cdot)$ associated to $\chi\in\{q^{v+\frac{t}{2}},q^{u-\frac{t}{2}}\}$ converge to those of $\mathfrak{p}_{t}^{d}(\cdot)$. When $v+\frac{t}{2}<0$ the atoms of $\hat{\pi}_{t}^{(N),d}(\cdot)$ are at $\hat{y}_{j}^{(N),v}(t)$ for $j\in[0,...,\lfloor-v-\frac{t}{2}\rfloor]$ and
$\lim_{N\to\infty}\hat{y}_{j}^{(N),v}(t)=x_{j}^{v}(t):=-4\left(v+j+\frac{t}{2}\right)^{2}.$
Similarly, when $u-\frac{t}{2}<0$ the atoms of $\hat{\pi}_{t}^{(N),d}(\cdot)$ are at $\hat{y}_{j}^{(N),u}(t)$ for $j\in[0,...,\lfloor u-\frac{t}{2}\rfloor]$ and
$\lim_{N\to\infty}\hat{y}_{j}^{(N),u}(t)=x_{j}^{u}(t):=-4\left(u+j+\frac{t}{2}\right)^{2}.$
However, it remains to show that their masses converge to the appropriate atomic masses for the continuous dual Hahn process. The arguments for the atoms arising when $02v+t<$ 0 are the same as those when the atoms arise from $02u-t<$ 0, and therefore we will only show the full argument in the former case.
Proof of Lemma 3.2 (2).
We recall the expression for $\hat{\pi}_{t}^{(N),d}(\hat{y}_{0}^{(N),v}(t))$,
$\hat{\pi}_{t}^{(N),d}(\hat{y}_{0}^{(N),v}(t))$ $=\frac{(q^{-2v-t},-q^{w+u},q^{w+r},-q^{r+u-t})_{\infty}}{(-q^{w-v},q^{u-v-t},-q^{r-v-t},q^{v+u+w+r})_{\infty}}.$
We can write the logarithm of this expression as
$\log{\left(N^{u+v}\hat{\pi}_{t}^{(N),d}(\hat{y}_{0}^{(N),v}(t))\right)}$ $=(u+v)\log{(N)}+\mathcal{A}^{+}_{N}[-2v-t,w+r]+\mathcal{A}^{-}_{N}[w+u,r+u-t]$
$-\mathcal{A}^{+}_{N}[u-v-t,u+v+w+r]-\mathcal{A}^{-}_{N}[w-v,r-v-t]$
$+E^{+}_{N}[-2v-t,w+r]+E^{-}_{N}[w+u,r+u]$
$-E^{+}_{N}[u-v-t,u+v+w+r]-E^{-}_{N}[w-u,r-v-t].$
The first five terms on the right side simplify to
$\log{(C_{u,v,w,r})}+\log{(\mathfrak{p}_{t}^{d}(x_{0}^{v}(t)))}=\log{\left(\frac{C_{u,v,w,r}\Gamma(u+v+2)\Gamma(u-v-t)}{\Gamma(-2v-t)}\mathfrak{p}_{t}^{d}(x_{0}^{v}(t))\right)}.$
To bound the error terms we simply note that all the arguments of the $E_{N}^{\pm}$ functions are constants, and thus we may apply the bound from Lemma 3.5 (2) directly. As a result, we see that
(3.7) $N^{u+v}\hat{\pi}_{t}^{(N),d}(\hat{y}_{0}^{(N),v}(t))=C_{u,v,w,r}\cdot\mathfrak{p}_{t}^{d}(x_{0}^{v}(t))\cdot e^{\textsc{Error}_{t}^{d}(x_{0}^{u}(t))},$
and for all $b\in(0,1)$, there exists a $C>0$ such that $|\textsc{Error}_{t}^{d}(x_{0}^{u}(t))|\leq CN^{-\frac{b}{2}}.$
The remaining atoms have weights $\hat{\pi}_{t}^{(N),d}(\hat{y}_{j}^{(N),v}(t))$,
$\hat{\pi}_{t}^{(N),d}(\hat{y}_{j}^{(N),v}(t))$ $=\hat{\pi}_{t}^{(N),d}(\hat{y}_{0}^{(N),v}(t))$
×$\times\frac{(q^{2v+t},-q^{v+w+t},q^{v+u},-q^{v+r})_{j}(1-q^{2v+2j+t})}{(q,-q^{1+v-w},q^{1+v-u+t},-q^{1+v-r+t})_{j}(1-q^{2v+t})(q^{w+r+u+v-1})^{j}}.$
showing a similar convergence result for the measure on these atoms reduces to showing a bound on the multiplicative factor. The following lemma follows from equations $(9.58)$ and $(9.59)$ in [CK24].
Lemma 3.8 (p. 58 [CK24]).
For all $a\in\mathbb{R}$ if we define functions $E_{N}^{1}(x),E_{N}^{2}(x)$ and $E_{N}^{3}(x)$ such that
$\frac{1-q^{a}}{1-q}=a\cdot e^{E_{N}^{1}(a)},$ $\frac{1+q^{a}}{1+q}=e^{E_{N}^{2}(a)},$ $q^{a}=e^{E^{3}_{N}(a)},$
then there exist constants $C_{a}>0$ such that
$|E_{N}^{1}(a)|,|E_{N}^{2}(a)|,|E_{N}^{3}(a)|\leq C_{a}N^{-\frac{1}{2}}.$
Applying this lemma, we can rewrite the multiplicative terms which depend on $j$ as
$\frac{(v+j+\frac{t}{2})[2v+t,v+u]_{j}}{(v+\frac{t}{2})j![1+v-u+t]_{j}}e^{E_{N}^{M_{j}}},$
and there exist $C\in\mathbb{R}_{>0}$ such that $|E_{N}^{M_{j}}|\leq CN^{-\frac{1}{2}}$. Therefore, applying the definition of $\mathfrak{p}_{t}^{d}(x_{j}^{v}(t))$ from 2.2, we see that
(3.8) $N^{u+v}\hat{\pi}_{t}^{v}(y_{j}^{(N),d})=C_{u,v,w,r}\cdot\mathfrak{p}_{t}^{d}(x_{j}^{v}(t)).$
The argument for the atoms arising from $u-\frac{t}{2}$ is identical, and will therefore not be repeated.
As a result of (3.7) and (3.8) and the analogous results for the atoms arising from $u-\frac{t}{2}$, we see that
$\lim_{N\to\infty}N^{u+v}\hat{\pi}_{t}^{v}(y_{j}^{(N),d})$ $=C_{u,v,w,r}\cdot\mathfrak{p}_{t}^{d}(x_{j}^{v}(t))$
$\lim_{N\to\infty}N^{u+v}\hat{\pi}_{t}^{u}(y_{j}^{(N),d})$ $=C_{u,v,w,r}\cdot\mathfrak{p}_{t}^{d}(x_{j}^{u}(t))$
which completes the proof of Lemma 3.2 (2). ∎
3.2. Proof of Lemma 3.3
Proof of Lemma 3.3 (1).
We begin by recalling the expression for $\hat{\pi}_{s,t;v,w,u,r}^{(N),c,c}(m,r)$:
(3.9) $\hat{\pi}_{s,t;v,w,u,r}^{(N),c,c}(m,r)=\frac{1}{2N}\pi_{q^{t},q^{s}}^{(N),c,c}\left(1-\frac{r}{2N},1-\frac{m}{2N}\right)\cdot\frac{\hat{\pi}_{t;v,w,u,r}^{(N),c}(r)}{\hat{\pi}_{s;v,w,u,r}^{(N),c}(m)}$
Therefore, we need to understand the asymptotics of the following Askey-Wilson measure expression.
$\sqrt{1-\frac{m}{4N}}\frac{1}{2N}\pi_{q^{t},q^{s}}\left(1-\frac{r}{2N},1-\frac{m}{2N}\right)=$
$\frac{\left(q,-q^{v+w+s},q^{t-s}\right)_{\infty}\bigg{|}\left(q^{v+\frac{t}{2}+i\frac{\sqrt{N}\theta_{r}}{2}},-q^{w+\frac{t}{2}+i\frac{\sqrt{N}\theta_{r}}{2}},q^{i\frac{\sqrt{N}\theta_{m}}{2}}\right)\bigg{|}^{2}}{4\pi N\sqrt{\frac{m}{N}}\left(-q^{v+w+t}\right)_{\infty}\bigg{|}\left(q^{v+\frac{s}{2}+i\frac{\sqrt{N}\theta_{m}}{2}},-q^{w+\frac{s}{2}+i\frac{\sqrt{N}\theta_{m}}{2}},q^{\frac{t-s}{2}+i\frac{\sqrt{N}(\theta_{r}+\theta_{m})}{2}},q^{\frac{t-s}{2}+i\frac{\sqrt{N}(-\theta_{r}+\theta_{m})}{2}}\right)_{\infty}\bigg{|}^{2}}.$
With a slight modification of the notation we used in Section 3.1, defined below, we can rewrite the logarithm of this equation.
(3.10) $\log{\left(\sqrt{1-\frac{m}{4N}}\frac{1}{2N}\pi_{q^{t},q^{s}}^{(N),c,c}\left(1-\frac{r}{2N},1-\frac{m}{2N}\right)\right)}=$
$-\log{\left(4\pi N\sqrt{\frac{m}{N}}\right)}+\mathcal{A}_{N}(\sqrt{m},\sqrt{r})+E_{N}^{\mathcal{A}}(m,r)+E_{N}^{\mathcal{E}}(m,r)$
In the expression above,
$\mathcal{A}_{N}(\sqrt{m},\sqrt{r})$ $:=\mathcal{A}_{N}^{+}\left[1,t-s,v+\frac{t}{2}+i\frac{\sqrt{r}}{2},v+\frac{t}{2}-i\frac{\sqrt{r}}{2},i\sqrt{m},-i\sqrt{m}\right]$
$+\mathcal{A}_{N}^{-}\left[v+w+s,w+\frac{t}{2}+i\frac{\sqrt{r}}{2},w+\frac{t}{2}-i\frac{\sqrt{r}}{2}\right]$
$-\mathcal{A}^{+}_{N}\left[v+\frac{s}{2}+i\frac{\sqrt{m}}{2},v+\frac{s}{2}-i\frac{\sqrt{m}}{2},\frac{t-s}{2}+i\frac{\sqrt{r}+\sqrt{m}}{2}\right]$
$-\mathcal{A}^{+}_{N}\left[\frac{t-s}{2}-i\frac{\sqrt{r}+\sqrt{m}}{2},\frac{t-s}{2}+i\frac{\sqrt{r}-\sqrt{m}}{2},\frac{t-s}{2}-i\frac{\sqrt{r}-\sqrt{m}}{2}\right]$
$-\mathcal{A}^{-}_{N}\left[v+w+t,w+\frac{s}{2}+i\frac{\sqrt{m}}{2},w+\frac{s}{2}-i\frac{\sqrt{m}}{2}\right],$
$E_{N}^{\mathcal{A}}(m,r)$ $:=\mathcal{A}_{N}(\sqrt{m}+E_{N}^{\theta}(m),\sqrt{r}+E_{N}^{\theta}(r))-\mathcal{A}_{N}(\sqrt{m},\sqrt{r}),$
$E^{\mathcal{E}}_{N}(m,r)$ $:=\mathcal{E}_{N}(\sqrt{m}+E^{\theta}_{N}(m),\sqrt{r}+E^{\theta}_{N}(r)).$
As before, $E_{N}^{\theta}(a)$ is defined by $\sqrt{N}\theta_{a}^{(N)}=\sqrt{a}+E_{N}^{\theta}(a).$ We can simplify the first two terms of (3.10) by expanding using the definitions of $\mathcal{A}^{\pm}$, (2.13), and then canceling like terms.
$-\log{\left(4\pi N\sqrt{\frac{r}{N}}\right)}+\mathcal{A}_{N}(\sqrt{m},\sqrt{r})=$
$\log{\left(\frac{\bigg{|}\Gamma\left(v+\frac{s}{2}+i\frac{\sqrt{m}}{2}\right)\Gamma\left(\frac{t-s}{2}+i\frac{\sqrt{r}+\sqrt{m}}{2}\right)\Gamma\left(\frac{t-s}{2}+i\frac{\sqrt{r}-\sqrt{m}}{2}\right)\bigg{|}^{2}}{8\pi\sqrt{r}\Gamma(t-s)\bigg{|}\Gamma\left(v+\frac{t+i\sqrt{r}}{2}\right)\Gamma\left(\frac{i\sqrt{m}}{2}\right)\bigg{|}^{2}}\right)}$
We also simplify the expression for the ratio of the one-point distributions which arises in (3.9).
$\log{\left(\frac{\sqrt{1-\frac{r}{4N}}\hat{\pi}_{t;v,w,u,r}^{(N),c}(r)}{\sqrt{1-\frac{m}{4N}}\hat{\pi}_{s;v,w,u,r}^{(N),c}(m)}\right)}=\log{\left(\frac{\mathfrak{p}_{t}^{c}(r)}{\mathfrak{p}_{s}^{c}(m)}\right)}+\textsc{Error}_{t}^{(N),c}(r)-\textsc{Error}_{s}^{(N),c}(m)$
using these expressions, and the definitions of the continuous dual Hahn process, 2.2, we can rewrite (3.9) as
$\log{\left(\sqrt{1-\frac{r}{4N}}\pi_{s,t}^{(N),c,c}(m,r)\right)}$ $=E_{N}^{\mathcal{A}}(m,r)+E_{N}^{\mathcal{E}}(m,r)+\textsc{Error}_{t}^{(N),c}(r)-\textsc{Error}_{s}^{(N),c}(m)$
$+\log{\left(\frac{\bigg{|}\Gamma\left(v+\frac{s}{2}+i\frac{\sqrt{m}}{2}\right)\Gamma\left(\frac{t-s}{2}+i\frac{\sqrt{r}+\sqrt{m}}{2}\right)\Gamma\left(\frac{t-s}{2}+i\frac{\sqrt{r}-\sqrt{m}}{2}\right)\bigg{|}^{2}\mathfrak{p}_{t}^{c}(r)}{8\pi\sqrt{r}\Gamma(t-s)\bigg{|}\Gamma\left(v+\frac{t}{2}+i\frac{\sqrt{r}}{2}\right)\Gamma\left(i\frac{\sqrt{m}}{2}\right)\bigg{|}^{2}\mathfrak{p}_{s}^{c}(m)}\right)}$
$=\log{\left(\mathfrak{p}_{s,t}^{c,c}(m,r)\right)}+E_{N}^{\mathcal{A}}(m,r)+E_{N}^{\mathcal{E}}(m,r)+\textsc{Error}_{t}^{(N),c}(r)-\textsc{Error}_{s}^{(N),c}(m)$
The last step is to bound the error terms. The bounds on $\textsc{Error}_{t}^{(N),c}(r)$ and $\textsc{Error}_{s}^{(N),c}(m)$ were already given in Section 3.1. The remaining error bound is stated in the following lemma.
Lemma 3.9.
For all $\eta>1$ there exists $N_{0}\in\mathbb{Z}_{\geq 1}$ and $C,\chi>0$ such that for all $N>N_{0}$ and $m,r\in\hat{S}^{(N),c}=[0,4N]$,
$|E_{N}^{\mathcal{A}}(m,r)|,|E_{N}^{\mathcal{E}}(m,r)|\leq CN^{-\chi}(1+\sqrt{m})^{\eta}+CN^{-\chi}(1+\sqrt{r})^{\eta}$
The proof of this lemma is the same as the proof of Lemma 3.6, the only change being the slight difference in definition of the error functions, so we will not repeat it. ∎
Proof of Lemma 3.3 (2).
We rewrite $\hat{\pi}_{s,t;v,w,u,r}^{(N),d,c}\left(\hat{y}_{j}^{(N),\chi}(s),r\right)$ through the same time reversal transformation:
(3.11) $\hat{\pi}_{s,t;v,w,u,r}^{(N),d,c}(\hat{y}_{j}^{(N),\chi}(s),r)=\pi_{q^{t},q^{s}}^{(N),c,d}\left(1-\frac{r}{2N},y_{j}^{(N),\chi}(s)\right)\cdot\frac{\hat{\pi}_{t;v,w,u,r}^{(N),c}(r)}{\hat{\pi}_{s}^{(N),d}(\hat{y}_{j}^{(N),\chi}(s))}$
We analyze the expression
$\pi_{q^{t},q^{s}}^{(N)}\left(1-\frac{r}{2N},V\right)=AW(V;q^{v+\frac{s}{2}},-q^{w+\frac{s}{2}},q^{\frac{t-s}{2}+i\frac{\sqrt{N}\theta_{r}}{2}},q^{\frac{t-s}{2}-i\frac{\sqrt{N}\theta_{r}}{2}})$
Atoms may arise when $02v+s<$ 0, with associated weights
(3.12) $\pi_{q^{t},q^{s}}^{(N),c,d}\left(1-\frac{r}{2N},y_{0}^{(N),v}(s)\right)$ $=\frac{(q^{-2v-s},-q^{w+\frac{t}{2}+i\frac{\sqrt{N}\theta_{r}}{2}},-q^{w+\frac{t}{2}-i\frac{\sqrt{N}\theta_{r}}{2}},q^{t-s})_{\infty}}{(-q^{w-v},q^{-v+\frac{t}{2}-s+i\frac{\sqrt{N}\theta_{r}}{2}},q^{-v+\frac{t}{2}-s-i\frac{\sqrt{N}\theta_{r}}{2}},-q^{v+t+w})_{\infty}},$
(3.13) $\pi_{q^{t},q^{s}}^{(N),c,d}\left(1-\frac{r}{2N},y_{j}^{(N),v}(s)\right)$ $=\pi_{q^{t},q^{s}}^{(N),c,d}\left(1-\frac{r}{2N},y_{0}^{(N),v}(s)\right)\times M_{j},$
(3.14) $M_{j}$ $=\frac{(q^{2v+s},-q^{v+w+s},q^{v+\frac{t}{2}+i\frac{\sqrt{N}\theta_{r}}{2}},-q^{v+\frac{t}{2}-i\frac{\sqrt{N}\theta_{r}}{2}})_{j}(1-q^{2v+2j+s})}{(q,-q^{1+v-w},q^{1+v+s-\frac{t}{2}-i\frac{\sqrt{N}\theta_{r}}{2}},q^{1+v+s-\frac{t}{2}+i\frac{\sqrt{N}\theta_{r}}{2}})_{j}(1-q^{2v+s})(-q^{v+w+t-1})^{j}}.$
To analyze the weights which arise from the atoms, we need to ensure that there are no instances of division by zero. Any division by zero in (3.12) arises when $-v+\frac{t}{2}-s\in\mathbb{Z}_{\leq 0}$ or, in (3.14), when $01+v+s-\frac{t}{2}+j-1=$ 0 for $j\in[[1,\lfloor-v-\frac{s}{2}\rfloor]]$, or when $02v+s=$ 0. The final case can never happen, since we assumed, in order to produce atoms, that $v+\frac{s}{2}<0$. Furthermore, since $s<t$, the other two scenarios are also ruled out. We see that
$\log{\left(\pi_{q^{t},q^{s}}^{(N),c,d}\left(1-\frac{r}{2N},y_{0}^{(N),v}(s)\right)\right)}=\mathcal{A}_{N}(\sqrt{r})+E_{N}^{\mathcal{A}}(r)+E_{N}^{\mathcal{E}}(r)$
Where, in this equation, we define $\mathcal{A}_{N}(a)$ for $a\in\mathbb{R}_{\geq 0}$ by
(3.15) $\mathcal{A}_{N}(\sqrt{r})$ $:=\mathcal{A}_{N}^{+}\left[-2v-s,t-s\right]+\mathcal{A}_{N}^{-}\left[b+\frac{t}{2}+i\frac{\sqrt{r}}{2},b+\frac{t}{2}-i\frac{\sqrt{r}}{2}\right]$
$-\mathcal{A}_{N}^{+}\left[-v+\frac{t}{2}-s+i\frac{\sqrt{r}}{2},-v+\frac{t}{2}-s-i\frac{\sqrt{r}}{2}\right]-\mathcal{A}_{N}^{-}\left[b-v,v+t+b\right]$
$=\log{\left(\frac{\bigg{|}\Gamma\left(\frac{t}{2}-s-v+\frac{i\sqrt{r}}{2}\right)\bigg{|}^{2}}{\Gamma(-2v-s)\Gamma(t-s)}\right)}$
$E_{N}^{\mathcal{A}}(r)$ $:=\mathcal{A}_{N}(\sqrt{r}+E^{\theta}(r))-\mathcal{A}_{N}(\sqrt{r}),$
$E_{N}^{\mathcal{E}}(r)$ $:=\mathcal{E}_{N}(\sqrt{r}+E_{N}^{\theta}(r)).$
We also need to address the ratio of $\hat{\pi}_{t}^{(N),c}(r)$ and $\hat{\pi}_{s}^{(N),d}(y_{j}^{(N),v}(s))$ which arose in (3.11).
$\log{\left(\frac{\sqrt{1-\frac{r}{4N}}\hat{\pi}_{t;v,w,u,r}^{(N),c}(r)}{\hat{\pi}_{s}^{(N),d}(\hat{y}_{j}^{(N),v}(s))}\right)}=\log{\left(\frac{\mathfrak{p}_{t}^{c}(y)}{\mathfrak{p}_{s}^{d}(x_{j}^{v})}\right)}+\textsc{Error}_{t}^{(N),c}(r)-\textsc{Error}_{s}^{(N),d}(x_{0}^{v}(s))$
Consequently, applying the definitions of the continuous dual Hahn process measures from 2.2, we find
(3.16) $\log{\left(\sqrt{1-\frac{r}{4N}}\hat{\pi}_{s,t;v,w,u,r}^{(N),d,c}(\hat{y}_{0}^{(N),v}(s),r)\right)}$ $=\log{\left(\frac{|\Gamma(\frac{t}{2}-s-v+\frac{i\sqrt{r}}{2})|^{2}\mathfrak{p}_{t}^{c}(r)}{\Gamma(-2v-s)\Gamma(t-s)\mathfrak{p}_{s}^{d}(x_{j}^{v})}\right)}+E_{N}^{\mathcal{A}}(r)$
$+E_{N}^{\mathcal{E}}(r)+\textsc{Error}_{t}^{(N),c}(r)-\textsc{Error}_{s}^{(N),d}(x_{0}^{v}(s))$
$=\log{\left(\mathfrak{p}_{s,t}^{d,c}(x_{0}^{v}(s),r)\right)}+E_{N}^{\mathcal{A}}(r)+E_{N}^{\mathcal{E}}(r)$
$+\textsc{Error}_{t}^{(N),c}(r)-\textsc{Error}_{s}^{(N),d}(x_{0}^{v}(s))$
To finish the argument, we need the following error bounds.
Lemma 3.10.
For $E_{N}^{\mathcal{A}}(r)$ and $E_{N}^{\mathcal{E}}(r)$ as defined by (3.15), for all $\eta>1$ there exists $N_{0}\in\mathbb{Z}_{\geq 1}$ and $C,\chi>0$ such that for all $N>N_{0}$ and $r\in\hat{S}^{(N),c}=[0,4N]$,
$|E_{N}^{\mathcal{A}}(r)|,|E_{N}^{\mathcal{E}}(r)|\leq CN^{-\chi}(1+\sqrt{r})^{\eta}$
The proof of this lemma is unchanged from the earlier bound on the error terms in Lemma 3.6, the only change being the modified definitions of $E_{N}^{\mathcal{A}}(r)$ and $E_{N}^{\mathcal{E}}(r)$, and therefore we will not repeat the proof. Combining the bounds from Lemma 3.10 and Lemma 3.5, we see that for all $\eta>1$, there exist $C,\chi>0$ such that the error in (3.16) is bounded by $CN^{-\chi}(1+\sqrt{r})^{\eta}$. This bound completes the argument in the case that $j=0$, and all that remains is to investigate the terms in $M_{j}$, defined in (3.14).
Lemma 3.11 (p. 64, [CK24]).
The multiplicative term $M_{j}$ can be written as
(3.17) $M_{j}=\frac{[2v+s,v+\frac{t}{2}+i\frac{\sqrt{r}}{2},v+\frac{t}{2}-i\frac{\sqrt{r}}{2}]_{j}(2v+2j+s)}{[1,1+v+s-\frac{t}{2}+i\frac{\sqrt{r}}{2},1+v+s-\frac{t}{2}-i\frac{\sqrt{r}}{2}]_{j}(2v+s)}(-1)^{j}e^{E_{N}^{M_{j}}(r)}$
Where, for all $\eta>1$ there exists $N_{0}\in\mathbb{Z}_{\geq 1}$ and $C,\chi>0$ such that for all $N>N_{0}$ and $r\in\hat{S}_{t}^{(N),c}=[0,4N]$,
$|E_{N}^{M_{j}}(r)|\leq CN^{-\chi}(1+\sqrt{r})^{\eta}$
Applying this lemma, we see that
$\log{\left(\sqrt{1-\frac{r}{4N}}\hat{\pi}_{s,t;v,w,u,r}^{(N),d,c}(\hat{y}_{j}^{(N),v}(s),r)\right)}=\log{\left(\mathfrak{p}_{s,t}^{d,c}(x_{j}^{v},r)\right)}+\overline{\textsc{Error}}_{s,t}^{(N),d,c}(x_{j}^{v}(s),r)$
With $\overline{\textsc{Error}}_{s,t}^{(N),d,c}(x_{j}^{v}(s),r):=\textsc{Error}_{s,t}^{(N),d,c}(x_{j}^{v}(s),r)+E_{N}^{M_{j}}$. By the above calculations, for all $\eta>1$, there exist $N_{0}\in\mathbb{Z}_{\geq 1}$ and $C,\chi>0$ such that for all $N>N_{0}$, $r\in\hat{S}^{(N),c}_{t}$,
$\overline{\textsc{Error}}_{s,t}^{(N),d,c}(x_{j}^{v}(s),r)\leq CN^{-\chi}(1+\sqrt{r})^{\eta}$
Finally, we can give the proof of Lemma 3.3 (3).
Proof of Lemma 3.3 (3).
Here $u-\frac{s}{2}<0$ and consequently, $v+\frac{s}{2}\geq 0$. We note that the Askey-Wilson measure associated to this case will be
$\pi_{q^{t},q^{s}}^{(N)}\left(1-\frac{r}{2N},V\right)=AW(V;q^{v+\frac{s}{2}},-q^{w+\frac{s}{2}},q^{\frac{t-s}{2}+i\frac{\sqrt{N}\theta_{r}}{2}},q^{\frac{t-s}{2}-i\frac{\sqrt{N}\theta_{r}}{2}})$
The only possibility is that we have atoms arising from $w+\frac{t}{2}<0$, but we have restricted our domain in time so that these atoms do not arise. ∎
3.3. Proof of Lemma 3.4
We begin with the proof of part (1).
Proof of Lemma 3.4 (1).
We know that the time-reversed transition measure can be expressed via the following transformation on the standard Askey-Wilson measure.
$\hat{\pi}_{s,t}^{(N),d,d}(x,y)=\pi_{q^{t},q^{s}}^{(N),d,d}\left(1-\frac{y}{2N},1-\frac{x}{2N}\right)\cdot\frac{\hat{\pi}_{t}^{(N),d}(y)}{\hat{\pi}_{s}^{(N),d}(x)}$
The relevant Askey-Wilson measure is
$\pi_{q^{t},q^{s}}^{(N)}\left(y_{j}^{(N),v}(t),V\right)$ $=AW\left(V;q^{v+\frac{s}{2}},-q^{w+\frac{s}{2}},q^{\frac{t}{2}+v+j},q^{\frac{t}{2}-s-v-j}\right)$
This measure will only have atoms arising from $v+\frac{s}{2}<0$ and $w+\frac{s}{2}<0$, however, we restrict the domain in $t$ so that the latter case never occurs. What remains are atoms arising from $v+\frac{s}{2}<0$, which have weights
$\pi_{q^{t},q^{s}}^{(N)}\left(y_{j}^{(N),v}(t),y_{0}^{(N),v}(s)\right)$ $=\frac{(q^{-2v-s},-q^{w+t+j+v},-q^{w-j-v},q^{t-s})_{\infty}}{(-q^{w-v},q^{t-s+j},q^{-2v-s-j},-q^{v+w+t})_{\infty}}$
$\pi_{q^{t},q^{s}}^{(N)}\left(y_{j}^{(N),v},y_{k}^{(N),v}\right)$ $=\pi_{q^{t},q^{s}}^{(N)}\left(y_{j}^{(N),v},y_{0}^{(N),v}\right)\times M_{k}$
$M_{k}$ $=\frac{(q^{2v+s},-q^{v+w+s},q^{2v+t+j},q^{-j})_{k}(1-q^{2v+2k+s})}{(q,-q^{1+v-w},q^{1+s-t-j},q^{1+2v+s+j})_{k}(1-q^{2v+s})(-q^{v+w+t-1})^{k}}$
Taking the logarithm, we write
$\log{\left(\pi_{q^{t},q^{s}}^{(N)}\left(y_{j}^{(N),v}(t),y_{0}^{(N),v}(s)\right)\right)}$ $=\mathcal{A}_{N}(j)+E_{N}^{\mathcal{A}}(j)+E_{N}^{\mathcal{E}}(j),$
where $\mathcal{A}_{N}(j)$ is defined
$\mathcal{A}_{N}(j)$ $:=\mathcal{A}^{+}_{N}[-2v-s,t-s]+\mathcal{A}^{-}[w+t+j+v,w-j-v]$
$-\mathcal{A}^{+}[t-s+j,-2v-s-j]-\mathcal{A}^{-}[w-v,v+w+t]$
$=\log{\left(\frac{\Gamma(t-s+j)\Gamma(-2v-s-j)}{\Gamma(-2v-s)\Gamma(t-s)}\right)}.$
As before, albeit with modified error functions $E_{N}^{\mathcal{A}}(j)$ and $E_{N}^{\mathcal{E}}(j)$, we have the following lemma.
Lemma 3.12.
There exists $N_{0}\in\mathbb{Z}_{\geq 1}$ and $C,\chi>0$ such that for all $N>N_{0}$ and $j\in[[0,\lfloor-u-\frac{t}{2}\rfloor]]$, $E_{N}^{\mathcal{A}}(j)$ and $E_{N}^{\mathcal{E}}(j)$ are bounded by
$|E_{N}^{\mathcal{A}}(j)|,|E_{N}^{\mathcal{E}}(j)|\leq CN^{-\chi}$
Using the lemma, we find that
$\pi_{q^{t},q^{s}}^{(N),d,d}\left(y_{j}^{(N),v}(t),y_{0}^{(N),v}(s)\right)=\frac{\Gamma(t-s+j)\Gamma(-2v-s-j)}{\Gamma(-2v-s)\Gamma(t-s)}e^{\textsc{Error}_{t,s}^{(N),d,d}(y_{j}^{(N),v}(t),y_{0}^{(N),v}(s))}$
and there exist $C,\chi>0$ and $N_{0}\in\mathbb{Z}_{\geq 1}$ so that for all $N>N_{0}$,
$|\textsc{Error}_{s,t}^{(N),d,d}(y_{j}^{(N),v}(t),y_{0}^{(N),v}(s))|\leq CN^{-\chi}$
Now, we just need to deal with the $M_{k}$ terms. We can apply a version of Lemma 3.11 to conclude that
$M_{k}=\frac{[2v+s,2v+t+j,-j]_{k}(2v+2k+s)}{[1,1+s-t-j,1+2v+s+j]_{k}(2v+s)}(-1)^{k}e^{\overline{\textsc{Error}}_{t,s}^{(N),d,d}(y_{j}^{(N),v}(t),y_{k}^{(N),v}(s))}$
and there exist $C,\chi>0$ and $N_{0}\in\mathbb{Z}_{\geq 1}$ such that for all $N>N_{0}$,
$|\overline{\textsc{Error}}_{t,s}^{(N),d,d}(y_{j}^{(N),v}(t),y_{k}^{(N),v}(s))|\leq CN^{-\chi}$
This finishes the proof of Lemma 3.4, part (1). ∎
We can also deal with the limit of the ration of $\hat{\pi}_{t}^{(N),d}(y)$ and $\hat{\pi}_{s}^{(N),d}(x)$ using our one-point computations.
$\lim_{N\to\infty}\frac{\hat{\pi}_{t}^{(N),d}(\hat{y}_{j}^{(N),v})}{\hat{\pi}_{s}^{(N),d}(\hat{y}_{j}^{(N),v})}$ $=\lim_{N\to\infty}\frac{N^{u+v}\hat{\pi}_{t}^{(N),d}(\hat{y}_{j}^{(N),v}(t))}{N^{u+v}\hat{\pi}_{s}^{(N),d}(\hat{y}_{k}^{(N),v}(s))}=\frac{\mathfrak{p}_{t}^{d}(x_{j}^{v}(t))}{\mathfrak{p}_{s}^{d}(x_{j}^{v}(s))}$
Consequently, we can prove Lemma 3.4 (1).
$\lim_{N\to\infty}\hat{\pi}_{s,t}^{(N),d,d}(\hat{y}_{j}^{(N),v},\hat{y}_{k}^{(N),v})=\mathfrak{p}_{s,t}(x_{k}^{v}(s),x_{j}^{v}(t))$
We now turn to Lemma 3.4 (2).
Proof of Lemma 3.4 (2).
We begin by considering the transition measures,
$\hat{\pi}_{s,t}^{(N),d,d}\left(\hat{y}_{j}^{(N),u}(s),\hat{y}_{k}^{(N),v}(t)\right)$ $=\pi_{q^{t},q^{s}}^{(N),d,d}\left(y_{k}^{(N),v}(t),y_{j}^{(N),u}(s)\right)\frac{\pi_{q^{t}}^{(N),d}\left(y_{k}^{(N),v}(t)\right)}{\pi_{q^{s}}^{(N),d}\left(y_{j}^{(N),u}(s)\right)}$
$\pi_{q^{t},q^{s}}^{(N),d,c}\left(y_{k}^{(N),v}(t),V\right)$ $=AW\left(V;q^{v+\frac{s}{2}},-q^{w+\frac{s}{2}},q^{\frac{t-s}{2}+\left(k+v+\frac{t}{2}\right)},q^{\frac{t-s}{2}-\left(k+v+\frac{t}{2}\right)}\right)$
which has no $u$ atoms. Likewise,
$\hat{\pi}_{s,t}^{(N),d,d}\left(\hat{y}_{j}^{(N),u}(s),\hat{y}_{k}^{(N),u}(t)\right)$ $=\pi_{q^{t},q^{s}}^{(N),d,d}\left(y_{k}^{(N),u}(t),y_{j}^{(N),u}(s)\right)\frac{\pi_{q^{t}}^{(N),d}\left(y_{k}^{(N),u}(t)\right)}{\pi_{q^{s}}^{(N),d}\left(y_{j}^{(N),u}(s)\right)}$
$\pi_{q^{t},q^{s}}^{(N),d,d}\left(y_{k}^{(N),u}(t),V\right)$ $=AW\left(V;q^{v+\frac{s}{2}},-q^{w+\frac{s}{2}},q^{\frac{t}{2}-s+u+k},q^{\frac{t}{2}-u-k}\right)$
which also has no atoms from $u-\frac{s}{2}<0$. ∎
Finally, we prove Lemma 3.4 (3).
Proof of Lemma 3.4 (3).
We write
$\frac{1}{2N}\pi_{q^{t},q^{s}}^{(N)}\left(y_{j}^{(N),v}(t),1-\frac{r}{2N}\right)$ $=\frac{1}{4\pi N\sqrt{\frac{r}{N}}\sqrt{1-\frac{r}{4N}}}\times$
$\frac{(q,-q^{v+w+s},q^{t+j+2v},q^{-j},-q^{t+w+j+v},-q^{w-j-v},q^{t-s})_{\infty}|(q^{i\sqrt{N}\theta_{r}})|^{2}}{(-q^{v+w+t})_{\infty}\bigg{|}\left(q^{v+\frac{s}{2}+i\frac{\sqrt{N}\theta_{r}}{2}},-q^{w+\frac{s}{2}+i\frac{\sqrt{N}\theta_{r}}{2}},q^{\frac{t-s}{2}+(j+v+\frac{t}{2})+i\frac{\sqrt{N}\theta_{r}}{2}},q^{\frac{t-s}{2}-(j+v+\frac{t}{2})+i\frac{\sqrt{N}\theta_{r}}{2}}\right)_{\infty}\bigg{|}^{2}}$
The $(q^{-j})_{\infty}$ term in the numerator sends the entire expression to 00. We can consider instead the transition from the atoms associated to $u$.
$\frac{1}{2N}\pi_{q^{t},q^{s}}^{(N)}\left(y_{j}^{(N),u}(t),1-\frac{r}{2N}\right)$ $=\frac{1}{4\pi N\sqrt{\frac{r}{N}}\sqrt{1-\frac{r}{4N}}}\times$
$\frac{(q,-q^{v+w+s},q^{j+v+u},q^{v-j-u+t},-q^{w+j+u},-q^{w-j-u+t},q^{t-s})_{\infty}|(q^{i\sqrt{N}\theta_{r}})|^{2}}{(-q^{v+w+t})_{\infty}\bigg{|}\left(q^{v+\frac{s}{2}+i\frac{\sqrt{N}\theta_{r}}{2}},-q^{w+\frac{s}{2}+i\frac{\sqrt{N}\theta_{r}}{2}},q^{\frac{t-s}{2}+(j+u-\frac{t}{2})+i\frac{\sqrt{N}\theta_{r}}{2}},q^{\frac{t-s}{2}-(j+u-\frac{t}{2})+i\frac{\sqrt{N}\theta_{r}}{2}}\right)_{\infty}\bigg{|}^{2}}$
With the same asymptotic analysis as in prior sections, we find that
$\sqrt{1-\frac{r}{4N}}\frac{1}{2N}\pi_{q^{t},q^{s}}^{(N)}\left(y_{j}^{(N),u}(t),1-\frac{r}{2N}\right)=$
$\frac{\big{|}\Gamma\left(v+\frac{s}{2}+i\frac{\sqrt{r}}{2}\right)\Gamma\left(j+u-\frac{s}{2}+i\frac{\sqrt{r}}{2}\right)\Gamma\left(t-\frac{s}{2}-j-u+i\frac{\sqrt{r}}{2}\right)\big{|}^{2}}{8\pi\sqrt{r}\Gamma\left(j+v+u\right)\Gamma\left(v-j-u+t\right)\Gamma\left(t-s\right)|\Gamma(i\sqrt{r})|^{2}}\cdot e^{\textsc{Error}_{t,s}^{(N)}(y^{(N),u}_{j}(t))}$
with the error term bounded above for all $r\in\hat{S}^{(N),c}=[0,4N]$ as
$|\textsc{Error}_{s,t}^{(N),d,c}|\leq CN^{-\chi}(1+\sqrt{r})^{\eta}.$
4. Proof of Proposition 3.1
In this section, we put the results from Section 3 together to prove 1.3 $(2)$. We begin by defining functions
$\mathcal{G}^{(N)}(\textbf{r};\textbf{c};\textbf{x})$ $:=\mathbf{1}_{\textbf{r}\in\mathbb{R}^{d+1}_{\leq 4N}}\cdot 2^{-N}\prod_{k=1}^{d+1}\left(\cosh{\left(\frac{s_{k}}{\sqrt{N}}\right)}+1-\frac{r_{k}}{2N}\right)^{N(x_{k}^{(N)}-x_{k-1}^{(N)})}.$
$\mathcal{G}(\textbf{r};\textbf{c};\textbf{x})$ $:=\exp{\left(\frac{1}{4}\sum_{k=1}^{d}(s_{k}^{2}-r_{k})(x_{k}-x_{k-1})\right)}$
We make these definitions in order to concisely rewrite the formula for the Laplace transform of the diffusively scaled height function (2.1) in terms of $\mathcal{G}^{(N)}(\mathbf{r};\mathbf{c};\mathbf{x})$ and $\mathcal{G}(\mathbf{r};\mathbf{c};\mathbf{x})$,
$\varphi^{(N)}_{u,v,w,r}(\textbf{c},\textbf{x})$ $\frac{\mathbb{E}\left[\prod_{k=1}^{d+1}\left(\cosh{\left(s_{k}^{(N)}\right)}+\mathbb{Y}_{e^{-2s_{k}^{(N)}}}\right)^{n_{k}-n_{k-1}}\right]}{\mathbb{E}\left[(1+\mathbb{Y}_{1})^{N}\right]},$
$=\frac{\mathbb{E}\left[\mathcal{G}^{(N)}\left((\hat{\mathbb{Y}}_{s_{1}}^{(N)},...,\hat{\mathbb{Y}}^{(N)}_{s_{d+1}});\textbf{c};\textbf{x}\right)\right]}{\mathbb{E}\left[\mathcal{G}^{(N)}\left((\hat{\mathbb{Y}}_{0}^{(N)},...,\hat{\mathbb{Y}}^{(N)}_{0});\mathbf{0};\textbf{x}\right)\right]}=\frac{\mathbb{E}\left[\mathcal{G}^{(N)}\left((\hat{\mathbb{Y}}_{s_{1}}^{(N)},...,\hat{\mathbb{Y}}^{(N)}_{s_{d+1}});c;X\right)\right]}{2^{-N}\mathbb{E}\left[\left(1-\frac{\hat{\mathbb{Y}}_{0}^{(N)}}{4N}\right)^{N}\right]}.$
$\varphi_{u,v}(\textbf{c};\textbf{x})$ $:=\frac{\mathbb{E}\left[\exp{\left(\frac{1}{4}\sum_{k=1}^{d}\left(s_{k}^{2}-\mathbb{T}_{s_{k}}\right)\left(x_{k}-x_{k-1}\right)\right)}\right]}{\mathbb{E}\left[\left(-\frac{1}{4}\mathbb{T}_{0}\right)^{N}\right]}$
$=\frac{\mathbb{E}\left[\mathcal{G}\left((\mathbb{T}_{s_{1}},...,\mathbb{T}_{s_{d}});\textbf{c};\textbf{x}\right)\right]}{\mathbb{E}\left[\mathcal{G}\left((\mathbb{T}_{0},...,\mathbb{T}_{0});\textbf{c};\textbf{x}\right)\right]}.$
In particular, we break this ratio up into its numerator and denominator, defining two associated functions
$\widetilde{\varphi}_{u,v,w,r}^{(N)}(\textbf{c};\textbf{x})$ $:=\int\mathcal{G}^{(N)}(\textbf{r};\textbf{c};\textbf{x})\prod_{k=1}^{d}\hat{\pi}^{(N)}_{s_{i+1},s_{i}}(r_{i+1},dr_{i})\cdot N^{u+v}\hat{\pi}^{(N)}_{s_{d+1}}(dr_{d+1})$
$=N^{u+v}\mathbb{E}\left[\mathcal{G}^{(N)}\left((\hat{\mathbb{Y}}_{s_{1}}^{(N)},...,\hat{\mathbb{Y}}^{(N)}_{s_{d+1}});\textbf{c};\textbf{x}\right)\right]$
$\widetilde{\varphi}_{u,v}(\textbf{c};\textbf{x})$ $:=\int\mathcal{G}(\textbf{r};\textbf{c};\textbf{x})\prod_{k=1}^{d}\mathfrak{p}_{s_{i+1},s_{i}}(r_{i+1},dr_{i})\cdot\mathfrak{p}_{s_{d+1}}(dr_{d+1})$
$=\mathbb{E}\left[\mathcal{G}\left((\mathbb{T}_{s_{1}},...,\mathbb{T}_{s_{d}});\textbf{c};\textbf{x}\right)\right]$
In this language, we may write
$\varphi^{(N)}_{u,v,w,r}(\textbf{c};\textbf{x})=\frac{\widetilde{\varphi}^{(N)}_{u,v,w,r}(\textbf{c};\textbf{x})}{\widetilde{\varphi}^{(N)}_{u,v,w,r}(\mathbf{0};\textbf{x})}.$
The proof of 1.3 $(2)$ will proceed by demonstrating that (Lemma 4.3)
$\lim_{N\to\infty}\widetilde{\varphi}_{u,v,w,r}^{(N)}(\textbf{c};\textbf{x})=C_{u,v,w,r}\widetilde{\varphi}_{u,v}(\textbf{c};\textbf{x}).$
We cite the following lemma from Corwin and Knizel [CK24].
Lemma 4.1 (Lemma 8.8 [CK24]).
For all $N\in\mathbb{N}$, the functions $\widetilde{\varphi}_{u,v}(\textbf{c},\textbf{x})$, $\textbf{c}\in(0,C_{u,r})^{d}$, and $\widetilde{\varphi}_{u,v}(\textbf{0},\textbf{x})$ take values in $(0,\infty)$.
The proof is identical to that in [CK24], so we only paraphrase the argument here. The first step is to apply the bound $\mathcal{G}(\textbf{r};\textbf{c};\textbf{x})\leq Ce^{-(1-x_{d})r_{d+1}/4}$ and then repeatedly apply identities in Lemma 2.4 to integrate over the variables $r_{i}$ to obtain
$\widetilde{\varphi}_{v,w,u,r}(\textbf{c};\textbf{x})\leq C\int_{\mathbb{R}}e^{-(1-x_{d})r}\mathfrak{p}_{0}(dr).$
Then, it is necessary to apply further bounds to the asymptotic behavior of $\mathfrak{p}_{0}^{c}$ to show that this integral is finite. Showing that $\widetilde{\varphi}_{u,v}(\textbf{c};\textbf{x})$ is positive involves noting that $\mathcal{G}(\textbf{r};\textbf{c};\textbf{x})\geq\mathcal{G}(\mathbf{2};\textbf{c};\textbf{x})$ for all $\mathbf{r}\in[1,2]^{d+1}$ and then using this bound to obtain the lower bound
$\widetilde{\varphi}_{u,v}(\mathbf{c};\mathbf{x})\geq\mathcal{G}(\mathbf{2};\mathbf{c};\mathbf{x})\int_{[1,2]^{d+1}}\prod_{i=1}^{d}\mathfrak{p}_{s_{i+1},s_{i}}^{c,c}(r_{i+1},r_{i})dr_{i}\cdot\mathfrak{p}_{s_{d+1}}^{c}(r_{d+1})dr_{d+1}$
and then completing the bound by using the explicit formulas for the marginal and transition measures to integrate this expression.
We will also use the following lemma, analogous to Lemma 8.2 in [CK24].
Lemma 4.2 (Lemma 8.2 [CK24]).
For every compact interval $I\subset\mathbb{R}$, there exists a constant $C>0$ such that for all $\textbf{c}\in\mathbb{R}^{d+1}$,
$\mathcal{G}^{(N)}(\textbf{r};\textbf{c};\textbf{x})\leq C\prod_{k=1}^{d+1}e^{-(x_{k}-x_{k-1})\frac{r_{k}}{4}+\frac{|r_{k}|}{2N}}$
where x and c are defined in (2.2). Furthermore, for all $\textbf{r}^{(N)}\in\mathbb{R}^{d+1}$ with $\lim_{N\to\infty}\textbf{r}^{(N)}=\textbf{r},$
$\lim_{N\to\infty}\mathcal{G}^{(N)}(\textbf{r}^{(N)};\textbf{c};\textbf{x})=\mathcal{G}(\textbf{r};\textbf{c};\textbf{x})$
The proof of this lemma is also unchanged from the proof of the analogous lemma in [CK24]. It follows from an involved sequence of applying elementary identities and approximations.
Finally, we will show the convergence of the sequence $\{\widetilde{\varphi}^{(N)}_{u,v,w,r}(\textbf{c};\textbf{x})\}_{N\in\mathbb{N}}$.
Lemma 4.3.
For $(q,A,B,C,D)$ and $h_{u,v,w,r}^{(N)}(x)$ satisfying the scaling conditions defined in 1.1, and for all $\textbf{c}\in(0,C_{u,r})^{d}$ and $\textbf{c}=\textbf{0}$,
$\lim_{N\to\infty}\widetilde{\varphi}_{u,v,w,r}^{(N)}(\textbf{c};\textbf{x})=C_{u,v,w,r}\widetilde{\varphi}_{u,v}(\textbf{c};\textbf{x}).$
The argument is unchanged from that given in Corwin and Knizel’s paper [CK24], albeit using the lemmas we established in Section 3 for general $w$ and $r$ and the scaling conditions in 1.1. For this reason, we only provide a summary of the argument here.
There are two main steps. The first is to show pointwise convergence of the integrands. This is accomplished by combining the results of Lemma 4.2 with those of Lemma 3.2, Lemma 3.3, and Lemma 3.4. Then we must show dominated convergence to get the full result. Using the error bounds in Lemma 3.2, Lemma 3.3, and Lemma 3.4, and the fact that $e^{-cr}e^{CN^{-\chi}(1+\sqrt{r})^{\eta}}\leq C^{\prime}e^{-c^{\prime}r}$ for $C^{\prime}>0$ sufficiently large and $c^{\prime}>0$ sufficiently small, and the bound that
$\mathcal{G}^{(N)}(\textbf{r};\textbf{c};\textbf{x})\leq Ce^{-c(r_{1}+...+r_{d+1}},$
which follows from Lemma 4.2, we obtain the dominating function
(4.1) $\bigg{|}\mathcal{G}^{(N)}(\textbf{r};\textbf{c};\textbf{x})\prod_{i=1}^{d}\hat{\pi}_{s_{i+1},s_{i}}^{(N),a_{i+1},a_{i}}(r_{i+1},r_{i})\cdot\hat{\pi}_{s_{d+1}}^{(N),a_{d+1}}(r_{d+1})\bigg{|},$
$\leq Ce^{-c(r_{1}+...+r_{d+1}}\prod_{i=1}^{d}\mathfrak{p}_{s_{i+1},s_{i}}^{a_{i+1},a_{i}}\cdot\mathfrak{p}_{s_{d+1}}^{a_{d+1}}(r_{d+1}).$
All that remains is to show that, for all $I\subset[[1,d+1]]$,
$\sum_{r_{I}}\int_{r_{J}}C^{\prime}e^{-c^{\prime}(r_{1}+...+r_{d+1})}\prod_{i=1}^{d}\mathfrak{p}_{s_{i+1},s_{i}}^{a_{i+1},a_{i}}(r_{i+1},r_{i})\cdot\mathfrak{p}_{s_{d+1}}^{a_{d+1}}(r_{d+1})<\infty,$
which is accomplished by casework on the structure of the set $I$.
5. Proof of Theorem 1.3
In this section, we discuss the tightness of the laws of $\{h_{u,v,w,r}^{(N)}(\cdot)-h_{u,v,w,r}^{(N)}(0)\}_{N\in\mathbb{N}}$, and then give a full proof of 1.3. A standard result in the study of stochastic processes (see, for instance, Karatzas and Shreve [KS91]) tells us that if, for all $0\leq x,y\leq 1$ there exist $\alpha,\beta,\nu,C\in\mathbb{R}_{>0}$ such that
(5.1) $\sup_{N\in\mathbb{N}}\mathbb{E}\left[|h_{u,v,w,r}^{(N)}(x)-h_{u,v,w,r}^{(N)}(y)|^{\alpha}\right]\leq C|x-y|^{1+\beta},$
and such that the process, evaluated at 00, has bounded supremum (which is immediate for our process since $h^{(N)}_{u,v,w,r}(0)-h^{(N)}_{u,v,w,r}(0)=0$), then the collection of laws of $\{h_{u,v,w,r}^{(N)}(\cdot)-h_{u,v,w,r}^{(N)}(0)\}_{N\in\mathbb{N}}$ is tight on $(C[0,1],\mathcal{B}(C[0,1]))$. Since $h_{u,v,w,r}^{(N)}(0)-h_{u,v,w,r}^{(N)}(0)=0$ for all $N\in\mathbb{N}$, the proof of tightness reduces to demonstrating the second bound in (5.1).
We will not prove the following lemma, as the argument is unchanged from that presented in [CK24]. The result relies on an attractive coupling of the open ASEP, which is also stated and proved in [CK24] Lemma 4.1. The argument proceeds by sandwiching the measure of $h^{(N)}_{u,v,w,r}(x)-h^{(N)}_{u,v,w,r}(y)$ between two stationary measures, $h_{u,-u,w,r}^{(N)}(x)-h_{u,-u,w,r}^{(N)}(0)$ and $h_{-v,v,w,r}^{(N)}(x)-h^{(N)}_{-v,v,w,r}(0)$ which are known (see [BW19], Remark 2.4) to be in product Bernoulli form.
Lemma 5.1 (Proposition 4.2 (4) [CK24]).
Consider the invariant measure for the open ASEP with parameters $A,B,C,D$ given by (1.3). Denote the associated diffusively scaled height function increment process by $\{h^{(N)}_{u,v,w,r}(x)-h^{(N)}_{u,v,w,r}(0)\}_{x\in[0,1]}$, as before. Then for all $u,v,w,r\in\mathbb{R}$, we have the following Hölder bound. For all $\theta\in(0,1/2)$ and every $n\in\mathbb{Z}_{\geq 1}$ there exists a constant $C(\theta,n,u,v)>0$ such that for every $x,y\in[0,1]$ and every $N\in\mathbb{Z}_{\geq 1}$
$\left\|h_{u,v,w,r}^{(N)}(x)-h_{u,v,w,r}^{(N)}(y)\right\|_{n}\leq C(\theta,n,u,v)|x-y|^{\theta},$
where $\|\cdot\|_{n}:=\mathbb{E}[(\cdot)^{n}]^{1/n}$.
This completes the proof of tightness for 1.3. It is important to note that, while Corwin and Knizel state Lemma 5.1 for the scaling conditions in 1.4, the proof relies on an attractive coupling of ASEP and a sandwiching argument for the diffusively scaled random walks $h_{u,-v,w,r}^{(N)}(x)$ and $h_{-v,v,w,r}^{(N)}(x)$ which is unchanged when we allow $w$ and $r$ to vary.
Finally, we can prove that our sequence of measures converges weakly to $H_{u,v}(\cdot)-H_{u,v}(0)$, completing the proof of 1.3.
Proof of 1.3.
We must show that, when $u+v>0$ and $w,r>0$, the finite dimensional distributions of $\{h_{u,v,w,r}^{(N)}(\cdot)-h_{u,v,w,r}^{(N)}(0)\}$ converge in distribution to those of $H_{u,v}(\cdot)-H_{u,v}(0)$. Using a standard result in the convergence of continuous stochastic processes (see, for example, Karatzas and Shreve [KS91], Ch. 2 Theorem 4.15), convergence in finite dimensional distributions with tightness of the sequence, from 1.3 $(1)$, implies that $\{h_{u,v,w,r}^{(N)}(\cdot)-h_{u,v,w,r}^{(N)}(0)\}\implies H_{u,v}(\cdot)-H_{u,v}(0)$ on $(C[0,1],\mathcal{B}(C[0,1])).$
For any $d\in\mathbb{Z}_{\geq 1}$ and $0<x_{1}<...<x_{d}\leq 1$ and $x^{(N)}=N^{-1}\lfloor Nx\rfloor$, we use $\mathbb{P}^{(N)}$ to denote the law of the vector $\left(h_{u,v,w,r}^{(N)}(x_{1}^{(N)})-h_{u,v,w,r}^{(N)}(0),...,h_{u,v,w,r}^{(N)}(x_{d}^{(N)})-h_{u,v,w,r}^{(N)}(0)\right)$. Consequently, using 1.3 $(1)$, we know that every subsequence of this collection has a convergent sub-subsequence. We know that the corresponding sequence of Laplace transforms converges to a unique limit, which is the Laplace transform of the open KPZ stationary measure on an open interval (3.1). Therefore, applying 2.6, we see that when $u+v>0$ and $w,r>0$, $\{h^{(N)}_{u,v,w,r}(\cdot)-h^{(N)}_{u,v,w,r}(0)\}\implies H_{u,v}(\cdot)-H_{u,v}(0).$